### v2

# On vector-valued functional equations with multiple recursive terms

Abstract
In this work, we study vector-valued functional equations with multiple recursive terms that arise naturally when we are dealing with vector-valued multiplicative Lindley-type recursions. Our work is strongly motivated by a wide range of semi-Markovian queueing, and vector-valued autoregressive processes. In all cases, we provide explicit expressions for the joint distribution in terms of Laplace-Stieltjes transforms/generating functions.
Keywords: Recursion; Laplace-Stieltjes transform; Generating functions; Markov-modulated arrivals and services; System of Wiener-Hopf type equations
1 Introduction
The primary aim of this work is to investigate vector-valued functional equations of the form
$\tilde{Z}(r,s,\eta)=G(r,s,\eta)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(r,\alpha_{i}(s),\eta)+K(r,s,\eta),$ (1)
for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$, where $\alpha_{i}(s)$, $i=1,\ldots,N$ are commutative contraction mappings, and $\tilde{Z}(r,s,\eta):=(Z_{1}(r,s,\eta),Z_{2}(r,s,\eta),\ldots,Z_{N}(r,s,\eta))^{T}$. Moreover, $G(r,s,\eta)$, $K(r,s,\eta)$ are given vector-valued functions and $\tilde{P}^{(i)}:=(\tilde{P}^{(i)})_{p,q}$, $i,p,q\in E$ is an $N\times N$ matrix, with the element $\tilde{P}^{(i)}_{i,i}=1$, and all the other elements $\tilde{P}^{(i)}_{p,q}=0$, $p,q\neq i$. Note that $\sum_{i=1}^{N}\tilde{P}^{(i)}=I$, that is the identity matrix. Such type of functional equations arise naturally when we are dealing with the time-dependent behaviour of vector-valued multiplicative Lindley-type recursions of a certain type. The corresponding stationary version results in the following vector-valued functional equation:
$\tilde{Z}(s)=H(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(\alpha_{i}(s))+\tilde{V}(s),\,Re(s)\geq 0,$ (2)
where $H(s)$, $V(s)$, are known matrix/vector-valued functions. A more general version of (2) is considered in Section 8. We also attempt to consider multidimensional versions such as
$\tilde{Z}(s,t)=K(s,t)+R(s,t)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(t,\alpha_{i}(t))+T(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(s,\alpha_{i}(t)),$ (3)
which arise in Section 7, and corresponds to the joint LST of a Markov-modulated tandem network of two queue with Levy input and consumption, where $K(s,t)$, $R(s,t)$, $T(s)$, are known matrix-valued functions. Moreover, we also focus on vector-valued reflected autoregressive processes that are described by recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where now $A$ is an $N\times N$ known matrix with elements $a_{i,j}\in[0,1)$, and our main concern is to provide expressions for the joint LST of the stationary vector $\tilde{Z}$, as a solution of a functional equation of the form (for $N=2$):
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j})-c_{2}(s_{1},s_{2})f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2})\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2})+c_{4}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}\lambda_{j},\sum_{j=1}^{2}a_{j,2}\lambda_{j}),\end{array}$ (4)
where $c_{i}(s_{1},s_{2})$, $i=1,2,3,4$ are known functions.
Our work is strongly motivated by a wide range of semi-Markovian queueing, and reflected autoregressive processes. In particular, (1) corresponds to a vector-valued analogue of equation (2) in [2]. Our primary aim in this work is to consider matrix generalizations of the seminal works in [2, 12]. Quite recently, the author in [20] studied vector-valued functional equations of the form
$\begin{array}[]{rl}\tilde{Z}(r,s,\eta)=&G(r,s,\eta)\tilde{Z}(r,\alpha(s),\eta)+K(r,s,\eta),\vskip 5.69054pt\\ \tilde{Z}(s)=&H(s)\tilde{Z}(a(s))+\tilde{V}(s),\end{array}$ (5)
where $\alpha(s)=as$, $a\in(0,1)$.
In this work, our aim is to extend the analysis in [20] to the case where the focus is on vector-valued functional equations of the form given in (5). Thus, we also extend the analysis in [2] to the matrix-valued framework. Functional equations of the form in (1), (5) arise naturally from the analysis of vector-valued recursions of the form
$\tilde{Z}_{n+1}=[R_{n}(X_{n})\tilde{Z}_{n}+\tilde{Y}_{n}(X_{n})-\tilde{B}_{n}(X_{n})]^{+},$ (6)
where $\{X_{n};n\in\mathbb{N}\}$ denotes an irreducible Markov chain with a finite state space $\{1,2,\ldots,N\}$, and vectors $\tilde{Z}_{n},\tilde{Y}_{n}(X_{n}),\tilde{B}_{n}(X_{n})\in\mathbb{R}^{n}$, and where $R_{n}(X_{n}),\tilde{Y}_{n}(X_{n}),\tilde{B}_{n}(X_{n})$ depend on the state of $X_{n}$, i.e., we consider Markov-modulated vector-valued recursions.
1.1 Related work
In the following, we provide a brief overview of the existed analytical results in the scalar case, since to our best knowledge, the only vector-valued version of (6) that has been treated analytically refers to the case where $R_{n}(X_{n})=a\in(0,1)$; see [20].
In [12], the authors considered the (scalar) recursion $Z_{n+1}=[aZ_{n}+S_{n}-A_{n+1}]^{+}$ (with $x]^{+}:=max\{x,0\}$), where $\{S_{n}-A_{n+1}\}_{n\in\mathbb{N}_{0}}$ forms a sequence of independent and identically distributed (i.i.d.) random variables and $a\in(0,1)$. The authors provided explicit results for the case where $\{S_{n}\}_{n\in\mathbb{N}_{0}}$ being a sequence of independent $exp(\lambda)$ distributed random variables, and $\{Y_{n}\}_{n\in\mathbb{N}_{0}}$ i.i.d., nonnegative and independent of $\{B_{n}\}_{n\in\mathbb{N}_{0}}$ with distribution function $F_{Y}(.)$ and Laplace–Stieltjes transform (LST) $\phi_{Y}(.)$. Note that in such a case, $Z_{n}$ could be interpreted as the workload in a queueing system just before the $n$th arrival, which adds $Y_{n}$ work, and makes obsolete a fixed fraction 1a1-a of the already present work. The case where $a=1$ corresponds to the classical Lindley recursion describing the waiting time of the classical G/G/1 queue, while the case where a=1a=-1 was investigated in [32]. Further progress has recently been made in [10] where the scalar autoregressive process described by the recursion $Z_{n+1}=[V_{n}Z_{n}+S_{n}-A_{n+1}]^{+}$ was investigated. In [11], the authors considered the case where $V_{n}W_{n}$ was replaced by $F(W_{n})$, where $\{F(t)\}$ is a Levy subordinator (recovering also the case in [12], where $F(t)=at$). Recently, in [8], the authors motivated by applications that arise in queueing and insurance risk models, they considered Lindley-type recursions where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}_{0}}$ obey a semi-linear dependence. These recursions can also be treated as of autoregressive type. Furthermore, the authors in [2] developed a method to study functional equations that arise in a wide range of queueing, autoregressive and branching processes. Finally, the author in [22], considered a generalized version of the model in [10], by assuming $V_{n}$ to take values in $(-\infty,1]$. In [3, eq. (9)], the authors considered a queueing system with two classes of impatient customers, leading to a specific example of (1) that was analyzed in detail; see also [24]. Quite recently, in [19] the authors generalized the work in [12] by considering, among others, non-trivial dependence structures among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$. We also mention the work in [9] where, among others, the authors attempted to provide analytical results for ASIP (asymmetric inclusion process) tandem queues with consumption, the analysis of which leads to a functional equation that is similar to those considered in reflected autoregressive processes.
A primary motivation for our work is related to recent results on vector-valued reflected autoregressive processes. Quite recently, in [20], the author investigated vector-valued reflected autoregressive processes, where the sequences $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ are governed by an irreducible background Markovian process with finite state space, i.e., he considered Markov-modulated reflected autoregressive processes. In particular, he assumed that each transition of the Markov chain generates a new interarrival time $A_{n+1}$ and its corresponding service time $S_{n}$, thus, considered the Markov-dependent version of the process analysed in [12]. Note that the specific case of $a=1$ in [20] corresponds to the waiting time in a single server queue with Markov-dependent interarrival and service times studied in [1]. Markov-dependent structure of the form considered in [1] has also been used in insurance mathematics; see [4]. The process analysed in [1] (i.e., for $a=1$) is a special case of the class of processes studied in [6], although in [1], all the results were given explicitly. The case where a=1a=-1 was investigated in [31] (see also [32, Chapter 5]) in the context of carousel models. In [20] the author focused on the case where $a\in(0,1)$. Moreover, contrary to the case considered in [1, 31], in which given the state of the Markov chain at times $n$, $n+1$, the distributions of $A_{n+1}$, $S_{n}$ are independent of one another for all $n$ (although their distributions depend on the state of the background Markov chain), he further considered the case where there is a dependence among $\{S_{n}\}_{n\in\mathbb{N}_{0}}$, $\{A_{n}\}_{n\in\mathbb{N}}$ based on Farlie-Gumbel-Morgenstern (FGM) copula. More precisely, $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ form a sequence of i.i.d. random vectors with a distribution function defined by FGM copula and dependent on the state of the underlying discrete time Markov chain. He also considered the case where $\{(S_{n},A_{n+1})\}_{n\in\mathbb{N}_{0}}$ have a bivariate matrix-exponential distribution, which is dependent on the state of the underlying discrete time Markov chain, as well as the case where there is a linear dependence among them. Special treatment was also given to the analysis of the case where the service times were dependent on the waiting time, as well as to the case where the server’s speed is workload proportional, i.e., a modulated shot-noise queue. The time-dependent analysis of a Markov-modulated reflected autoregressive process was also investigated.
1.2 Our contribution
Our primary goal is to explore a class of vector-valued stochastic recursions (6), in which some independence assumptions are lifted and for which, nevertheless, a detailed exact analysis can be provided. More precisely, by coping with the transient analysis of recursions (6), we result in a vector-valued functional equations of the form in (1), thus, extending considerably the work in [20], where vector-valued functional equations of the form in (5) were investigated, as well as the seminal works in [2, 12], where scalar functional equations of the form in (1) and (5) were studied, respectively. To solve this equation we make use of Liouville’s theorem [30] and Wiener-Hopf boundary value theory [15], although the vector-valued form requires additional technicalities. We also considered the stationary analysis of a simplified version of (6), that lead to a functional equation of type given in (2). We show that the solution of such type of equation does not necessarily requires the use of Liouville’s theorem and the reduction to a Wiener-Hopf boundary value problem. To cope with this kind of functional equations, we use as a vehicle Markov-dependent reflected autoregressive processes and other related models, for which the analysis results in functional equations of the form in (1), (2). We also go one step further and studied and consider extensions, under which given the state of the background Markov chain there is additional dependence structure based on the Farlie-Gumbel-Morgenstern (FGM) copula, i.e., $\tilde{Y}_{n}(X_{n})$, $\tilde{B}_{n}(X_{n})$ are not conditionally independent, but instead, they are dependent based on the FGM copula. In this work, we have mainly restricted ourselves to the case where $\alpha_{i}(s)$ are commutative contraction mappings. In particular, we focus on the case where $\alpha_{i}(s):=a_{i}s$. However, in section 6 we deal with a commutative mapping that is not a contraction, i.e., $\alpha_{i}(s)=s+\mu_{i}c$, $i=1,\ldots,N$.
We further consider a Markov-modulated fluid flow model with consumption, by generalising existed work in fluid queues by incorporating the concept of consumption. We also cope with a modulated ASIP tandem queue with Lévy input and consumption by extending the analysis of the model in [9] from the scalar to the Markov-modulated case, which in turn leads to a vector valued two-dimensional functional equation. Note that in such a case, we are dealing with the LST of the joint buffer content. In Section 8, we consider an even more general version of (2), related to the generating function of stationary queue-length distribution of a Markov-modulated $M/G/1$ queue with a general impatience scheme. This model is described by an integer-valued Markov-modulated reflected autoregressive process.
Finally, in Section 9, we focus on the stationary analysis of a vector-valued autoregressive process (VAR(1)) that is described by recursions of the form $\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n}]^{+}$, where now $A$ is an $N\times N$ known matrix. Our aim is to derive the LST of the stationary joint distribution of the process under study by solving a vector-valued functional equation of the form in (4). Despite the fact that the form of (4) is quite complicated, and the arguments of each component of the LST are linear combinations of $s_{i}$, $i=1,\ldots,N$, we succeed to solve it in case where the argument of $f(.)$ in the right hand side of (4) is an $N$-dimensional contraction mapping.
1.3 Organization of the paper
The remainder of the paper is organized as follows. In Section 2 we focus on the time-dependent analysis of a Markov-modulated reflected autoregressive process, where, among others, the autoregressive parameter depends on the state of an exogenous finite state irreducible Markov process. The analysis results on a functional equation of the form in (1), where $\alpha_{i}(s)=a_{i}s$, $a_{i}\in(0,1)$, $i=1,\ldots,N$. The transient analysis of a Markov-modulated fluid queue with consumption is treated in Section 3. In Section 4, we investigate the stationary behaviour of a special case of the model presented in Section 2, as well as an extension that incorporates dependencies based on the FGM copula. Section 5 is devoted to the stationary analysis of a modulated shot-noise queue, while in Section 6 we focus on the stationary behaviour of a modulated Markovian queue with dependencies among service time and waiting time. The stationary behaviour of a Markov-modulated ASIP tandem queue with consumption is investigated in Section 7. An integer vector-valued reflected autoregressive process is investigated in Section 8. Section 9 refers to the stationary analysis of reflected VAR(1) processes. Conclusions and some topics for further research are presented in Section 10.
2 The Markov modulated $M/G/1$-type reflected autoregressive process
Consider an M/G/1-type queue with Markov modulated arrivals and services. Let $\{X(t);t\geq 0\}$ be the background process that dictates the arrivals and services. $\{X(t);t\geq 0\}$ is a Markov chain on $E=\{1,2,\ldots,N\}$ with infinitesimal generator $Q=(q_{i,j})_{i,j\in E}$, and denote its stationary distribution by $\widehat{\pi}=(\pi_{1},\ldots,\pi_{N})$, i.e., $\widehat{\pi}Q=0$, and $\widehat{\pi}\mathbf{1}=1$, where $\mathbf{1}$ is the $N\times 1$ column vector with all components equal to 1. Denote also by $I$ the $N\times N$ identity matrix, and by $M^{T}$ the transpose of the matrix $M$.
Customers arrive at time epochs $T_{1},T_{2},\ldots$, $T_{1}=0$, and service times are denoted by $S_{1},S_{2},\ldots$. If $X(t)=i$, arrivals occur according to a Poisson process with rate $\lambda_{i}>0$ and an arriving customer has a service time $S_{i}$, with cumulative distribution function (cdf) $B_{i}(.)$, Laplace-Stieltjes transform (LST) $\beta_{i}^{*}(s):=\int_{0}^{\infty}e^{-st}dB_{i}(t)$, $\bar{b}_{i}=-\beta_{i}^{*\prime}(0)$, $i=1,\ldots,N$, and $B^{*}(s):=diag(\beta_{1}^{*}(s),\ldots,\beta_{N}^{*}(s))$. We assume that given the state of the background process $\{X(t);t\geq 0\}$, $S_{1},S_{2},\ldots$ are independent and independent of the arrival process. Let $A_{n}=T_{n}-T_{n-1}$, $n=2,3,\ldots$, and $Y_{n}=X_{T_{n}}$, $n=1,2,\ldots$. We assume that such an arrival makes obsolete a fixed fraction $1-a_{i}$, given that $Y_{n}=i$, $i\in E$. Denote now by $R_{n}(X_{n})$ a random variable with support $\{a_{1},\ldots,a_{N}\}$, with $a_{i}\in(0,1)$, and such that $P(R_{n}(X_{n})=a_{i}|X_{n}=i)=1$.
Our focus is on the derivation of the transient distribution (in terms of Laplace-Stieltjes transform) of the process $\{(W_{n},T_{n});n=1,2,\ldots\}$ in which $\{T_{n};n=1,2,\ldots\}$ is an increasing time sequence generated by the input process, and $W_{n}$ denotes the workload in the system just before the $n$th customer arrival that takes place at $T_{n}$. This model generalizes the work in [20, Section 5], in which the author considered the case where the autoregressive parameter is independent of the state of the background state.
Assume that $W_{1}=w$ and let for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$,
$Z_{j}^{w}(r,s,\eta):=\sum_{n=1}^{\infty}r^{n}E(e^{-sW_{n}-\eta T_{n}}1_{\{Y_{n}=j\}})|W_{1}=w),\,j=1,\ldots,N,$
and $\tilde{Z}^{w}(r,s,\eta)=(Z_{1}^{w}(r,s,\eta),\ldots,Z_{N}^{w}(r,s,\eta))^{T}$. Let also
$A_{i,j}(s)=E(e^{-sA_{n}}1_{\{Y_{n}=j\}}|Y_{n-1}=i),\,i,j\in E,$
and denote by $A(s)$ the $N\times N$ matrix with elements $A_{i,j}(s)$, $Re(s)\geq 0$. Following the lines in [27, Lemma 2.1], we have that $A(s)=M^{-1}(s)\Lambda$ where $\Lambda:=diag(\lambda_{1},\ldots,\lambda_{N})$ and $M(s)=sI+\Lambda-Q$. Let also
$V_{j}^{w}(r,s,\eta):=\sum_{n=1}^{\infty}r^{n+1}E((1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w),\,j=1,\ldots,N,$
with $\tilde{V}^{w}(r,s,\eta):=(V_{1}^{w}(r,s,\eta),\ldots,V_{N}^{w}(r,s,\eta))^{T}$, and let $p_{j}:=P(X_{0}=j)$, $j=1,\ldots,N$ with $\widehat{p}:=(p_{1},\ldots,p_{N})^{T}$.
Theorem 1
For $Re(s)=0$, $Re(\eta)\geq 0$, $|r|<1$,
$Z^{w}_{j}(r,s,\eta)-rp_{j}e^{-sw}=r\sum_{i=1}^{N}Z^{w}_{i}(r,a_{i}s,\eta)\beta_{i}^{*}(s)A_{i,j}(\eta-s)+V_{j}^{w}(r,s,\eta),\,j=1,2,\ldots,N,$ (7)
or equivalently, in matrix notation,
$\tilde{Z}^{w}(r,s,\eta)-r\Lambda(M^{T}(\eta-s))^{-1}B^{*}(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}^{w}(r,a_{i}s,\eta)=re^{-sw}\widehat{p}+\tilde{V}^{w}(r,s,\eta),$ (8)
where $\tilde{P}^{(i)}:=(\tilde{P}^{(i)})_{p,q}$, $i,p,q\in E$ is an $N\times N$ matrix, with the element $\tilde{P}^{(i)}_{i,i}=1$, and all the other elements $\tilde{P}^{(i)}_{p,q}=0$, $p,q\neq i$. Note that $\sum_{i=1}^{N}\tilde{P}^{(i)}=I$.
Proof. Using the identity (where $x^{+}:=max(x,0)$, $x^{-}:=min(x,0)$),
$e^{-sx^{+}}+e^{-sx^{-}}=e^{-sx}+1,$ (9)
we have for $Re(s)=0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{l}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{+}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =E((e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]}+1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{0}=w)+E((1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$
Note that
$\begin{array}[]{r}E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta(A_{n+1}+T_{n})}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =E(e^{-sR_{n}W_{n}-\eta T_{n}}e^{-sS_{n}-(\eta-s)A_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =\sum_{i=1}^{N}E(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\end{array}$
Thus, for $Re(s)=0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{rl}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=&\sum_{i=1}^{N}E(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\vskip 5.69054pt\\ &+E((1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$ (10)
Multiplying (10) by $r^{n+1}$ and taking the sum of $n=1$ to infinity gives:
$Z^{w}_{j}(r,s,\eta)-rE(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=j\}}|W_{1}=w)=r\sum_{i=1}^{N}Z^{w}_{i}(r,a_{i}s,\eta)\beta_{i}^{*}(s)A_{i,j}(\eta-s)+V_{j}^{w}(r,s,\eta).$
Note that $T_{1}=0$ and,
$E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=j\}}|W_{1}=w)=E(e^{-sw}1_{\{X_{0}=j\}})=e^{-sw}P(X_{0}=j)=e^{-sw}p_{j}.$
Substituting back in (10), we obtain the system of Wiener-Hopf equations (7), which in matrix notation, is equivalent to (8).
The following lemma is taken from [20]; see also [27].
Lemma 2
  1. 1.
    The $N$ eigenvalues, say $\nu_{i}$, $i=1\ldots,N$, of $\Lambda-Q^{T}$ all lie in $Re(s)>0$.
  2. 2.
    The $N$ zeros of $det((\eta-s)I+\Lambda-Q^{T})=0$ for $Re(\eta)\geq 0$, say $\mu_{i}(\eta)$, $i=1,\ldots,N$, are all in $Re(s)>0$ (i.e., For $Re(s)=0$, $Re(\eta)\geq 0$, $det((\eta-s)I+\Lambda-Q^{T})\neq 0$), and such that $\mu_{i}(\eta)=\nu_{i}+\eta$, $i=1,\ldots,N$.
Proof. Note that $\Lambda-Q^{T}-sI:=S(s)+R(s)$, where $R(s)=diag(\lambda_{1}+q_{1}-s,\ldots,\lambda_{N}+q_{N}-s)$, and $S(s):=(s_{i,j}(s))_{i,j=1,\ldots,N}:=\begin{pmatrix}0&q_{2,1}&\ldots&q_{N,1}\\ q_{1,2}&0&\ldots&q_{N,2}\\ \vdots&\vdots&\vdots\\ q_{1,N}&q_{2,N}&\ldots&0\\ \end{pmatrix}$. Moreover,
$|\lambda_{i}+q_{i}-s|\geq\lambda_{i}+q_{i}-|s|>q_{i}=|-q_{i,i}|=\sum_{j\neq i}|q_{i,j}|=\sum_{j=1}^{N}|s_{i,j}(s)|.$
Thus, from [16, Theorem 1, Appendix 2], for $Re(s)>0$, the number of zeros of $det(\Lambda-Q^{T}-sI)$ are equal to the number of zeros of $det(R(s))=\prod_{i=1}^{N}(\lambda_{i}+q_{i}-s)$. Assume now that all $\nu_{i}$ are distinct.
Similarly, for $Re(\eta)\geq 0$, $Re(s)=0$, $det(\Lambda-Q^{T}+(\eta-s)I)\neq 0$ (thus, the inverse of $M^{T}(\eta-s)$ exists), and all the zeros of $det(\Lambda-Q^{T}+(\eta-s)I)$ lie in $Re(s)>0$. It is easy to realize that these zeros, say $\mu_{i}(\eta)$, are such that $\mu_{i}(\eta)=\nu_{i}+\eta$, $i=1,\ldots,N$, where $\nu_{i}$ the eigenvalues of $\Lambda-Q^{T}$.
Thus,
$(M^{T}(\eta-s))^{-1}=\frac{1}{\prod_{i=1}^{N}(s-\mu_{i}(\eta))}L(\eta-s),$
where $L(\eta-s):=cof(M^{T}(\eta-s))$ is the cofactor matrix of $M^{T}(\eta-s)$.
Remark 3
Note that $(M^{T}(\eta-s))^{-1}$ can also be written as (see [27, equation (3.16)]):
$(M^{T}(\eta-s))^{-1}=R\ diag(\frac{1}{\mu_{1}(\eta)-s},\ldots,\frac{1}{\mu_{N}(\eta)-s})R^{-1},$
where $R$ is the matrix with the $i$th column, say $R_{i}$, $i=1,\ldots,N$, being the right eigenvector of $\Lambda-Q^{T}$ corresponding to the eigenvalue $\nu_{i}$.
Let $G_{i,j}(\eta,s):=(\Lambda L(\eta-s)B^{*}(s))_{i,j}$, $i,j=1,\ldots,N$. Then, (8) can be written as:
$\prod_{i=1}^{N}(s-\mu_{i}(\eta))[Z_{j}^{w}(r,s,\eta)-re^{-sw}p_{j}]-r\sum_{i=1}^{N}Z_{i}^{w}(r,a_{i}s,\eta)G_{i,j}(\eta,s)=\prod_{i=1}^{N}(s-\mu_{i}(\eta))V_{j}^{w}(r,s,\eta).$ (11)
Note that for $|r|<1$, $Re(\eta)\geq 0$,
  • The left-hand side of (11) is analytic in $Re(s)>0$, continuous in $Re(s)\geq 0$, and it is also bounded.
  • The right-hand side of (11) is analytic in $Re(s)<0$, continuous in $Re(s)\leq 0$, and it is also bounded.
Thus, Liouville’s theorem [30, Th. 2.52], implies that, in their respective half planes, both the left and the right hand side of (11) can be rewritten as a polynomial of at most $N$th degree in $s$, dependent of $r$, $\eta$, for large $s$: For $|r|<1$, $Re(\eta)\geq 0$, $Re(s)\geq 0$:
$\prod_{i=1}^{N}(s-\mu_{i}(\eta))[Z_{j}^{w}(r,s,\eta)-re^{-sw}p_{j}]-r\sum_{i=1}^{N}Z_{i}^{w}(r,a_{i}s,\eta)G_{i,j}(\eta,s)=\sum_{i=0}^{N}s^{i}C_{i,j}^{w}(r,\eta).$ (12)
Note that for $s=0$, (12) (having in mind that $G_{i,j}(\eta,0)=A_{i,j}(\eta)$) yields
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[Z_{j}^{w}(r,0,\eta)-rp_{j}]-r\sum_{i=1}^{N}Z_{i}^{w}(r,0,\eta)A_{i,j}(\eta)=C^{w}_{0,j}(r,\eta).$
However, from (11), for $s=0$,
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[Z_{j}^{w}(r,0,\eta)-rp_{j}]-r\sum_{i=1}^{N}Z_{i}^{w}(r,0,\eta)A_{i,j}(\eta)=0,$
since $V^{w}_{j}(r,0,\eta)=0$, $j=1,\ldots,N$. Thus, $C_{0,j}^{w}(r,\eta)=0$, so that $C_{0}^{w}(r,\eta)=(C_{0,1}^{w}(r,\eta),\ldots,C_{0,N}^{w}(r,\eta))^{T}=\mathbf{0}$.
Remark 4
This result may also be derived as follows. Note that, (8) is now written for $|r|<1$, $Re(s)=0$, $Re(\eta)\geq 0$ as:
$\prod_{i=1}^{N}(s-\mu_{i}(\eta))[\tilde{Z}^{w}(r,s,\eta)-re^{-sw}\widehat{p}]-r\Lambda L(\eta-s)B^{*}(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}^{w}(r,a_{i}s,\eta)=\prod_{i=1}^{N}(s-\mu_{i}(\eta))\tilde{V}^{w}(r,s,\eta).$ (13)
Note that for $|r|<1$, $Re(\eta)\geq 0$,
  • The left-hand side of (13) is analytic in $Re(s)>0$, continuous in $Re(s)\geq 0$, and it is also bounded.
  • The right-hand side of (13) is analytic in $Re(s)<0$, continuous in $Re(s)\leq 0$, and it is also bounded since $|E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}-\eta T_{n+1}}1(Y_{n+1}=j)|W_{1}=w)|\leq 1$, $Re(s)\leq 0$, $Re(\eta)\geq 0$.
Thus, by analytic continuation we can define an entire function such that it is equal to the left-hand side of (13) for $Re(s)\geq 0$, and equal to the right-hand side of (13) for $Re(s)\leq 0$ (with $|r|<1$, $Re(\eta)\geq 0$). Hence, by (a variant of) Liouville’s theorem [26] (for vector-valued functions; see also [28, p. 81, Theorem 3.32] or [18, p. 232, Theorem 9.11.1], or [5, p. 113, Theorem 3.12]) behaves as a polynomial of at most $N$th degree in $s$. Thus, for $Re(\phi)\geq 0$,
$\prod_{i=1}^{N}(s-\mu_{i}(\eta))[\tilde{Z}^{w}(r,s,\eta)-re^{-sw}\widehat{p}]-r\Lambda L(\eta-s)B^{*}(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}^{w}(r,a_{i}s,\eta)=\sum_{i=0}^{N}s^{i}C^{w}_{i}(r,\eta),$ (14)
where $C^{w}_{i}(r,\eta):=(C^{w}_{i,1}(r,\eta),\ldots,C^{w}_{i,N}(r,\eta))^{T}$, $i=0,1,\ldots,N$, column vectors still have to be determined.
Note that for $s=0$, (14) yields
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[(I-r\Lambda(M^{T}(\eta))^{-1})\tilde{Z}^{w}(r,0,\eta)-r\widehat{p}]=C^{w}_{0}(r,\eta).$ (15)
However, from (13), for $s=0$,
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[(I-r\Lambda(M^{T}(\eta))^{-1})\tilde{Z}^{w}(r,0,\eta)-r\widehat{p}]=\mathbf{0},$ (16)
since $\tilde{V}(r,0,\eta)=\mathbf{0}$, where $\mathbf{0}$, is $N\times 1$ column vector with all components equal to 0. Thus, $C^{w}_{0}(r,\eta)=\mathbf{0}$.
Note that (14) has the same form as eq. (50) in [2], although it is in matrix form. Moreover, (13) is the matrix analogue of [2, equation (48)], having, for $|r|<1$, $Re(\eta)\geq 0$, the form (see also [2, equation (2)])
$\mathbf{f}(r,s,\eta)=\mathbf{g}(r,s,\eta)\sum_{i=1}^{N}\tilde{P}^{(i)}\mathbf{f}(r,\alpha_{i}s,\eta)+\mathbf{K}(r,s,\eta),$ (17)
with $\mathbf{g}(r,s,\eta):=rA^{T}(\eta-s)B^{*}(s)$, $\mathbf{K}(r,s,\eta):=re^{-sw}\widehat{p}+\tilde{V}^{w}(r,s,\eta)$. Note that $\mathbf{g}(r,0,\eta)=r\Lambda(M^{T}(\eta))^{-1}=rA^{T}(\eta)$, $\mathbf{K}(r,0,\eta)=r\widehat{p}\neq\mathbf{0}$. Note that for $|r|<1$, $Re(\eta)\geq 0$ $|\mathbf{g}_{i,j}(r,0,\eta)|<1$, $|\mathbf{K}_{i,j}(r,0,\eta)|<1$, $i,j\in E$. Moreover, if for a matrix $M=(M_{i,j})_{i,j\in E}$, $M||=max_{1\leq i\leq N}\sum_{j=1}^{N}|M_{i,j}|$, then $\mathbf{g}(r,0,\eta)||<1$, $\mathbf{K}(r,0,\eta)||<1$.
It seems that a matrix generalization of [2, Theorem 2] can be applied to solve such kind of functional equations.
Therefore, for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$,
$\tilde{Z}^{w}(r,s,\eta)=rK^{w}(s,\eta)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}^{w}(r,a_{i}s,\eta)+L^{w}(r,s,\eta),$ (18)
where
$\begin{array}[]{rl}K^{w}(s,\eta):=&\Lambda(M^{T}(\eta-s))^{-1}B^{*}(s)=rA^{T}(\eta-s)B^{*}(s),\vskip 5.69054pt\\ L^{w}(r,s,\eta):=&re^{-sw}\widehat{p}+\frac{1}{\prod_{i=1}^{N}(s-\mu_{i}(\eta))}\sum_{i=1}^{N}s^{i}C_{i}^{w}(r,\eta).\end{array}$
Iterating (18) $n$ times yields
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(r,s,\eta)\tilde{Z}^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta),\end{array}$ (19)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)K^{w}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta)\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(r,s,\eta):=I$, and $F_{i_{1},\ldots,i_{N}}(r,s,\eta)=O$ (that is, the zero matrix) if one of the indices is equal to 1-1.
The next step is to investigate the convergence as $n\to\infty$. Note that $\tilde{P}^{(k)}||_{1}=1$, $k=1,\ldots,N$. Moreover,
[…]
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)||\leq|r|^{n+1}\binom{i_{1}+\ldots+i_{N}}{i_{1},\ldots,i_{N}}C(\eta)(1+\epsilon)^{n+1}.$ (21)
Note also that $\lim_{n\to\infty}\tilde{Z}^{w}(r,a^{n}s,\eta)=\tilde{Z}^{w}(r,0,\eta)$ satisfies for $|r|<1$, $Re(\eta)\geq 0$:
$(I-r\Lambda(M^{T}(\eta))^{-1})\tilde{Z}^{w}(r,0,\eta)=r\widehat{p}\Leftrightarrow(I-rA^{T}(\eta))\tilde{Z}^{w}(r,0,\eta)=r\widehat{p}\Leftrightarrow\tilde{Z}^{w}(r,0,\eta)=r(I-rA^{T}(\eta))^{-1}\widehat{p}$,
since the eigenvalues of $rA^{T}(\eta)$ are all strictly less than one when $Re(\eta)\geq 0$. So, as $n\to\infty$, by using (21), the second term of the right-hand side of (19) vanishes. Therefore, (19) becomes
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{n=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=n}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta).\end{array}$ (22)
Note that in (22), there are still $N^{2}$ unknowns to be determined, i.e., the terms $C_{i,j}^{w}(r,\eta)$, $i,j=1,\ldots,N$. These terms can be derived by applying the following steps:
  • Substituting $s=\mu_{k}(\eta)$, $k=1,\ldots,N$ in (12) yield an $N^{2}\times N^{2}$ system of equations for the elements of $C^{w}_{j}(r,\eta)=(C^{w}_{j,1}(r,\eta),\ldots,C^{w}_{j,N}(r,\eta))^{T}$, $j=1,\ldots,N$:
    $r\sum_{i=1}^{N}Z_{i}^{w}(r,a_{i}\mu_{k}(\eta),\eta)G_{i,j}(\eta,a_{i}\mu_{k}(\eta))=\sum_{i=1}^{N}\mu_{k}^{i}(\eta)C^{w}_{i,j}(r,\eta),\,j=1,\ldots,N,$ (23)
    where $Z_{i}^{w}(r,a_{i}\mu_{k}(\eta),\eta)$, $i=1,\ldots,N$ be the $i$th element of $\tilde{Z}^{w}(r,a_{i}\mu_{k}(\eta),\eta)$.
  • Substitute $s=\mu_{k}(\eta)$, $k=1,\ldots,N$ in (22) to obtain expressions for $Z_{i}^{w}(r,a_{i}\mu_{k}(\eta),\eta)$, $i=1,\ldots,N$.
  • Substituting the resulting expressions of $Z_{i}^{w}(r,a_{i}\mu_{k}(\eta),\eta)$, $i=1,\ldots,N$ in (23) we obtain an $N^{2}\times N^{2}$ system of equations for the unknowns $C_{i,j}^{w}(r,\eta)$, $i,j=1,\ldots,N$.
3 A Markovian fluid flow model with consumption
Consider a general stochastic fluid model with a single infinite capacity buffer where the buffer content $X(t)$ increases continuously. An external background process $J(t)$, $t\geq 0$ with a finite state space $E=\{1,2,\ldots,N\}$ affects the rate of the buffer increase. In particular, when $J(t)=i$, $i\in E$, the buffer content increases continuously with rate $r_{i}\in(-\infty,+\infty)-\{0\}$ (for convenience we assume that $r_{i}\neq 0$, $i\in E$). When $J(t)=i$ the process remains for an exponential amount of time with rate $q_{i}$, and then may move to any state $j$. At transition epochs, a fraction $1-a_{i}$ of the buffer content is instantaneously removed (i.e., is consumed). Let $T_{0}=0$ and $T_{1},T_{2},\ldots$ be the transition epochs of the process $\{J(t);t\geq 0\}$ with $T_{1}>0$. Let also $W_{0}=v$ and for $n=1,2,\ldots$, $W_{n}=X(T_{n})$ and $Y_{n}=J(T_{n}^{-})$ is the state of $\{J(t);t\geq 0\}$ just before transition epochs.
Let $A_{n}=T_{n}-T_{n-1}$ be the inter-jump time, $n=1,2,\ldots$, and define $R_{n}=\sum_{k=1}^{N}r_{k}1_{\{Z_{n}=k\}}$, $V_{n}=\sum_{k=1}^{N}a_{k}1_{\{Z_{n}=k\}}$. We then have the recursion:
$W_{n+1}=[V_{n}W_{n}+R_{n+1}A_{n+1}]^{+}.$ (24)
Define for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{rl}Z_{i,j}(r,s,\eta,v)=&\sum_{n=1}^{\infty}r^{n}E(e^{-sW_{n}-\eta T_{n}}1_{\{Y_{n}=j\}}|Y_{1}=i,W_{0}=v),\\ Z_{i}^{0}(s,\eta,v)=&E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v),\end{array}$
and for $Re(s)\leq 0$,
$V_{i,j}(r,s,\eta,v)=\sum_{n=1}^{\infty}r^{n+1}E((1-e^{-s[W_{n}-R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n}=j\}}|Y_{1}=i,W_{0}=v),$
and for $Re(s)=0$,
$G_{i,j}(s,\eta)=E(e^{-(sr_{j}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{n}=i).$
Let $\mathbf{Z}(r,s,\eta,v)$, $\mathbf{G}(s,\eta)$, $\mathbf{V}(r,s,\eta,v)$ be $N\times N$ matrices with elements $Z_{i,j}(r,s,\eta,v)$, $G_{i,j}(s,\eta)$ and $V_{i,j}(r,s,\eta,v)$, respectively. Let also $\mathbf{Z}^{0}(s,\eta,v)=diag(Z_{1}^{0}(s,\eta,v),\ldots,Z_{N}^{0}(s,\eta,v))$. Note that
$\begin{array}[]{rl}Z_{i}^{0}(s,\eta,v)=&E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v)\\ =&\int_{0}^{\infty}e^{-s[v+r_{i}u]^{+}-\eta u}q_{i}e^{-q_{i}u}du.\end{array}$
For $r_{i}>0$,
$v+r_{i}u]^{+}=\left\{\begin{array}[]{ll}v+ur_{i},&u\geq-v/r_{i}\Leftrightarrow u\geq 0,\\ 0,&u\leq 0,\end{array}\right$.
and for $r_{i}<0$,
$v+r_{i}u]^{+}=\left\{\begin{array}[]{ll}v+ur_{i},&u\leq-v/r_{i},\\ 0,&u\geq-v/r_{i}.\end{array}\right$.
Simple calculations imply that
$Z_{i}^{0}(s,\eta,v)=\left\{\begin{array}[]{ll}e^{-sv}\frac{q_{i}}{q_{i}+sr_{i}+\eta},&r_{i}>0\\ \frac{q_{i}}{q_{i}+sr_{i}+\eta}[e^{-sv}-e^{-(\eta+q_{i})v/r_{i}}]+\frac{q_{i}e^{-(\eta+q_{i})v/r_{i}}}{\eta+q_{i}},&r_{i}<0.\end{array}\right$.
Note also that,
$\mathbf{G}(s,\eta)=P\mathbf{q}(\mathbf{q}+s\mathbf{r}+\eta I)^{-1},$
where $P$ is the one-step probability matrix of $\{Y_{n};n\in\mathbb{N}\}$, and $\mathbf{q}:=diag(q_{1},q_{2},\ldots,q_{N})$, $\mathbf{r}:=diag(r_{1},r_{2},\ldots,r_{N})$.
Then, we have the following result.
Theorem 5
For $Re(s)=0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{r}[Z_{i,j}(r,s,\eta,v)-\delta_{i,j}Z_{i}^{0}(s,\eta,v)](r_{j}s+\eta+q_{j})-r\sum_{k=1}^{N}Z_{i,k}(r,a_{k}s,\eta,v)p_{k,j}q_{j}\vskip 5.69054pt\\ =(r_{j}s+\eta+q_{j})V_{i,j}(r,s,\eta,v).\end{array}$ (25)
Proof. Using the identity (9), we have for $Re(s)=0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{l}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)=E(e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{+}-\eta(T_{n}-A_{n+1})}1_{\{Y_{n+1}=j\}})|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =E(e^{-sV_{n}W_{n}-\eta T_{n}-(sR_{n+1}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}})|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}E(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,V_{0}=v)E(e^{-s(R_{n+1}+\eta)A_{n+!}}1_{\{Y_{n+1}=j\}})|Y_{n}=k)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}E(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,V_{0}=v)G_{k,j}(s,\eta)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v).\end{array}$
Multiplying with $r^{n+1}$ and adding for all $n$ we obtain
$Z_{i,j}(r,s,\eta,v)-rE(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=j\}}|Y_{1}=i,V_{0}=v)=r\sum_{k=1}^{N}Z_{i,k}(r,a_{k}s,\eta,v)G_{k,j}(s,\eta)+V_{i,j}(r,s,\eta,v).$
Simple computations yield (25).
Note that for $|r|<1$, $Re(\eta)\geq 0$,
  • The left-hand side of (25) is analytic in $Re(s)>0$, continuous in $Re(s)\geq 0$, and is also bounded.
  • The right-hand side of (25) is analytic in $Re(s)<0$, continuous in $Re(s)\leq 0$, and is also bounded.
Liouville’s theorem [30] states that for $Re(s)\geq 0$,
$Z_{i,j}(r,s,\eta,v)-\delta_{i,j}Z_{i}^{0}(s,\eta,v)](r_{j}s+\eta+q_{j})-r\sum_{k=1}^{N}Z_{i,k}(r,a_{k}s,\eta,v)p_{k,j}q_{j}=c_{i,j}^{(0)}(r,\eta,v)+sc_{i,j}^{(1)}(r,\eta,v),$ (26)
and for $Re(s)\leq 0$,
$(r_{j}s+\eta+q_{j})V_{i,j}(r,s,\eta,v)=c_{i,j}^{(0)}(r,\eta,v)+sc_{i,j}^{(1)}(r,\eta,v).$ (27)
For $s=0$, (27) yields $c_{i,j}^{(0)}(r,\eta,v)=0$.
Let $R^{+}=\{j\in E:r_{j}>0\}$, $R^{-}=\{j\in E:r_{j}<0\}$. Note that for $j\in R^{+}$, $r_{j}s+\eta+q_{j}=0$ vanishes at $s:=s_{j}^{+}=-\frac{\eta+q_{j}}{r_{j}}<0$, while for $j\in R^{-}$, $r_{j}s+\eta+q_{j}=0$ vanishes at $s:=s_{j}^{-}=-\frac{\eta+q_{j}}{r_{j}}>0$. Then, using (27), for $s=s_{j}^{+}$, $j\in R^{+}$
$(r_{j}s_{j}^{+}+\eta+q_{j})V_{i,j}(r,s_{j}^{+},\eta,v)=s_{j}^{+}c_{i,j}^{(1)}(r,\eta,v),$
thus, $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{+}$. Setting $s=s_{j}^{-}$, $j\in R^{+}$ in (26) yields,
$r\sum_{k=1}^{N}Z_{i,k}(r,a_{k}s_{j}^{-},\eta,v)p_{k,j}q_{j}=s_{j}^{-}c_{i,j}^{(1)}(r,\eta,v).$ (28)
Equation (28) provides a system of equations for $c_{i,j}^{(1)}(r,\eta,v)$, $i\in E$, $j\in R^{-}$. However, we need an expression for the $Z_{i,k}(r,a_{k}s_{j}^{-},\eta,v)$, $i,k\in E$, in the left-hand side of (28). In a matrix terms, (26) is written as
$\mathbf{Z}(r,s,\eta,v)=r\sum_{k=1}^{N}\mathbf{Z}(r,a_{k}s,\eta,v)\tilde{P}^{(k)}\mathbf{G}(s,\eta)+\mathbf{L}(r,s,\eta,v),$ (29)
where
$\mathbf{L}(r,s,\eta,v)=r\mathbf{Z}^{0}(s,\eta,v)+s\mathbf{C}^{(1)}(r,\eta,v)[s\mathbf{r}+\eta I+\mathbf{q}]^{-1},$
where $\mathbf{C}^{(1)}(r,\eta,v)$ an $N\times N$ matrix with $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{+}$, and $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{-}$ will be found by (28).
Iterating (29) $n$ times yields
$\begin{array}[]{rl}\mathbf{Z}(r,s,\eta,v)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}\mathbf{F}_{i_{1},\ldots,i_{M}}(r,s,\eta,v)\mathbf{L}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}\mathbf{F}_{i_{1},\ldots,i_{N}}(r,s,\eta,v)\mathbf{Z}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v),\end{array}$ (30)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta,v)$ are recursively defined by
$\mathbf{F}_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}\mathbf{F}_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)\tilde{P}^{(k)}\mathbf{G}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta),$
[…]
$\mathbf{Z}(r,s,\eta,v)=\sum_{k=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=k}\mathbf{F}_{i_{1},\ldots,i_{M}}(r,s,\eta,v)\mathbf{L}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v).$ (31)
The expression (31) can be used in (28) to obtain the unknown terms $c_{i,j}^{(1)}(r,\eta,v)$, $i\in E$, $j\in R^{-}$.
4 The Markov dependent case
In the following, we cope with generalizing the work in [20, Section 2], by further assuming that the autoregressive parameter depends also on the state of a background Markov chain. In particular, we focus on the limiting counterpart of a special case of the model considered in Section 2; see also Remark 6.
Consider a FIFO single-server queue, and let $T_{n}$ $n$th arrival to the system with $T_{1}=0$. Define also $A_{n}=T_{n}-T_{n-1}$, $n=2,3,\ldots$, i.e., is the time between the $n$th and (n1)(n-1)th arrival. Let $S_{n}$ be the service time of the $n$th arrival, $n\geq 1$. We assume that the inter-arrival and service times are regulated by an irreducible discrete-time Markov chain $\{Y_{n},n\geq 0\}$ with state space $E=\{1,2,...,N\}$ and transition probability matrix $P:=(p_{i,j})_{i,j\in E}$. Let $\tilde{\pi}:=(\pi_{1},\ldots,\pi_{N})^{T}$ be the stationary distribution of $\{Y_{n};n\geq 0\}$. Let $W_{n}$ the workload in the system just before the $n$th customer arrival. Such an arrival adds $S_{n}$ work but makes obsolete a fixed fraction $1-a_{i}$ of the work that is already present in the system, given that $Y_{n}=i\in E$. Denote also by $Z_{i}^{n}(s):=E(e^{-sW_{n}}1_{\{Y_{n}=i\}})$, $Re(s)\geq 0$, $i=1,\ldots,N$, $n\geq 0$, and assuming the limit exists, define $Z_{i}(s)=\lim_{n\to\infty}Z_{i}^{n}(s)$, $i=1,\ldots,N$. Let also $\tilde{Z}(s)=(Z_{1}(s),\ldots,Z_{N}(s))^{T}$.
The sequences $\{A_{n}\}_{n\in\mathbb{N}}$ and $\{S_{n}\}_{n\in\mathbb{N}_{0}}$ are autocorrelated as well as cross-correlated. Assume that for $n\geq 0$, $x,y\geq 0$, $i,j=1,\ldots,N$:
$\begin{array}[]{l}P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i,A_{2},\ldots,A_{n},S_{1},\ldots,S_{n-1},Y_{1},\ldots,Y_{n-1})\vskip 5.69054pt\\ =P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i)=B_{i}(y)p_{i,j}(1-e^{-\lambda_{j}x}):=p_{i,j}B_{i}(y)G_{A,j}(x),\end{array}$ (32)
where $B_{i}(.)$, $G_{A,j}(.)$ denote the distribution functions of service and interarrival times, given $Y_{n}=i$, $Y_{n+1}=j$, respectively. Note that $A_{n+1}$, $S_{n}$, $Y_{n+1}$ are independent of the past given $Y_{n}$, and $A_{n+1}$, $S_{n}$ are conditionally independent given $Y_{n}$, $Y_{n+1}$. Let also $B^{*}(s)=diag(\beta_{1}^{*}(s),\ldots,\beta_{N}^{*}(s))$, where $\beta_{i}^{*}(s):=\int_{0}^{\infty}e^{-sy}dB_{i}(y)$, $L(s):=diag(\frac{\lambda_{1}}{\lambda_{1}-s},\ldots,\frac{\lambda_{N}}{\lambda_{N}-s})$, $\Lambda=diag(\lambda_{1},\ldots,\lambda_{N})$.
Remark 6
Note that following the notation in the previous subsection, we have $A_{i,j}(s):=E(e^{-sA_{n}}1(Y_{n}=j)|Y_{n-1}=i)=p_{i,j}\frac{\lambda_{j}}{\lambda_{j}+s}$.
Remark 7
Moreover, an extension to the case where $A_{n+1}|Y_{n+1}=j$ is of phase-type, e.g., a mixed Erlang distribution with cdf
$\begin{array}[]{c}G_{A,j}(x):=\sum_{m=1}^{M}q_{m}(1-e^{-\lambda_{j}x}\sum_{l=0}^{m-1}\frac{(\lambda_{j}x)^{l}}{l!}),\,x\geq 0,\end{array}$
can be handled at a cost of a more complicated expression.
Theorem 8
The transforms $Z_{j}(s)$, $j=1,\ldots,N$ satisfy the system
$Z_{j}(s)=\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(s)Z_{i}(a_{i}s)-\frac{s}{\lambda_{j}-s}v_{j},$ (33)
where $v_{j}:=\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(\lambda_{j})Z_{i}(a_{i}\lambda_{j})$, $j=1,\ldots,N$. Equivalently, in matrix notation, the transform vector $\tilde{Z}(s)$ satisfies
$\tilde{Z}(s)=H(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(a_{i}s)+\tilde{V}(s),$ (34)
where $\tilde{P}^{(i)}:=(\tilde{P}^{(i)})_{p,q}$, $i,p,q\in E$ is an $N\times N$ matrix, with the $(i,i)$-element $\tilde{P}^{(i)}_{i,i}=1$, and all the other elements $\tilde{P}^{(i)}_{p,q}=0$, $p,q\neq i$. Note that $\sum_{i=1}^{N}\tilde{P}^{(i)}=I$. Moreover, $H(s)=L(s)P^{T}B^{*}(s)$, $\tilde{V}(s):=s(I-L(s))\tilde{v}$, $\tilde{v}:=(v_{1},\ldots,v_{N})^{T}$.
Proof. From the recursion $W_{n+1}=[a_{i}W_{n}+B_{n}-J_{n}]^{+}$ (given $Y_{n}=i$) we obtain the following equation for the transforms $Z_{j}^{n+1}(s)$, $j=1,\ldots,N$:
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}})=\sum_{i=1}^{N}P(Y_{n}=i)E(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)E(e^{-s[a_{i}W_{n}+B_{n}-J_{n}]^{+}}1_{\{Y_{n+1}=j\}}|Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E(e^{-s[a_{i}W_{n}+B_{n}-J_{n}]^{+}}|Y_{n+1}=j,Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[E\left(\int_{0}^{a_{i}W_{n}+B_{n}}e^{-s(a_{i}W_{n}+B_{n}-y)}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)+E\left(\int_{a_{i}W_{n}+B_{n}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\lambda_{j}e^{-s(a_{i}W_{n}+B_{n})}\int_{0}^{a_{i}W_{n}+B_{n}}e^{-(\lambda_{j}-s)y}dy+\int_{a_{i}W_{n}+B_{n}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}}{\lambda_{j}-s}e^{-s(a_{i}W_{n}+B_{n})}(1-e^{-(\lambda_{j}-s)(a_{i}W_{n}+B_{n})})+e^{-\lambda_{j}(a_{i}W_{n}+B_{n})}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}e^{-s(a_{i}W_{n}+B_{n})}-se^{-\lambda_{j}(a_{i}W_{n}+B_{n})}}{\lambda_{j}-s}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[\frac{\lambda_{j}}{\lambda_{j}-s}Z_{i}^{n}(a_{i}s)\beta_{i}^{*}(s)-\frac{s}{\lambda_{j}-s}Z_{i}^{n}(a_{i}\lambda_{j})\beta_{i}^{*}(\lambda_{j})\right]\vskip 5.69054pt\\ =&\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}Z_{i}^{n}(a_{i}s)\beta_{i}^{*}(s)-\frac{s}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}Z_{i}^{n}(a_{i}\lambda_{j})\beta_{i}^{*}(\lambda_{j}).\end{array}$
Letting $n\to\infty$ so that $Z_{j}^{n}(s)$ tends to $Z_{j}(s)$ we get (33). Writing the resulting equations in matrix form we get (34).
Remark 9
It is easy to realize that (34) refers to the matrix analogue of equation (2) in [2]. Moreover, $\tilde{V}(0)=\mathbf{0}$, so that $\mathbf{1}\tilde{V}(0)=0$ and $H(0)=P^{T}$, so that by denoting by $H_{j}(0)$ the $j$th column of matrix $H(0)$, we will have $\mathbf{1}H_{j}(0)=\sum_{k=1}^{N}p_{j,k}=1$, $j=1,\ldots,N$, where $\mathbf{1}$ the $1\times N$ row vector of ones. So a matrix analogue of [2, Theorem 2] should give the solution to matrix functional equations of the form as given in (34).
Iterating $n$ times (34) yields,
$\tilde{Z}(s)=\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(s)V(a_{i_{1},\ldots,i_{N}}(s))+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(s)\tilde{Z}(a_{i_{1},\ldots,i_{N}}(s)),$ (35)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(s)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)H(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(s):=I$, and $F_{i_{1},\ldots,i_{N}}(s)=O$ (i.e., the zero matrix) if one of the indices equals 1-1.
Note that as $n\to\infty$, $Z_{j}(a_{i_{1},\ldots,i_{N}}(s))\to Z_{j}(0)=E(1_{\{Z=j\}})=P(Z=j)=\pi_{j}$, so that $\tilde{Z}(0)=\tilde{\pi}$. Alternatively, by letting $s=0$ in (34) and having in mind that $\tilde{V}(0)=\mathbf{0}$, and $H(0)=P^{T}$, (34) yields $\tilde{Z}(0)=P^{T}\tilde{Z}(0)$, thus, $\tilde{Z}(0)=\tilde{\pi}$.
Note that for $i_{1}+\ldots+i_{N}=n$, $|a_{i_{1},\ldots,i_{N}}(s)|\leq\kappa^{n}|s|$, where $\kappa=max(a_{1},\ldots,a_{N})$. Then,
[…]
where $j^{*}\in E$ is such that $|Z_{j^{*}}(a_{i_{1},\ldots,i_{N}}(s))-\pi_{j^{*}}|\geq|Z_{j}(a_{i_{1},\ldots,i_{N}}(s))-\pi_{j}|$, $\forall j\in E$.
Rewrite (35) as follows
[…]
Note that the third term in the right hand side of (36) converges to zero as $n\to\infty$. Remind that $H(0)=P^{T}$. Thus, since for $i_{1}+\ldots+i_{N}=n$, $|a_{i_{1},\ldots,i_{N}}(s)|\leq\kappa^{n}|s|$, we have that $H(a_{i_{1},\ldots,i_{N}}(s))$ with $i_{1}+\ldots+i_{N}=n$ is close to $H(0)=P^{T}$. In other words, $H(a_{i_{1},\ldots,i_{N}}(s))-P^{T}||_{1}<\epsilon<\kappa^{-1}-1$, where $A||_{1}=max_{1\leq j\leq N}(\sum_{i=1}^{N}|a_{i,j}|)$ (the maximum absolute column sum). So there is a constant $C$ such that for $i_{1}+\ldots+i_{N}=n$, $F_{i_{1},\ldots,i_{N}}(s)||_{1}\leq\binom{i_{1}+\ldots+i_{N}}{i_{1},\ldots,i_{N}}C(1+\epsilon)^{n}$. Therefore, each element in the second term in (36) is bounded by $C(1+\epsilon)^{n+1}\kappa^{n+1}|s|E(W1_{\{Y=j^{*}\}})$, which tends to zero as $n\to\infty$. The following theorem gives the main result.
Theorem 10
For $\tilde{V}(0)=\mathbf{0}$,
[…]
It remains to obtain the vector $\tilde{v}$. This is given by the following proposition.
Proposition 11
The vector $\tilde{v}$ is the unique solution of the following system of equations:
$v_{j}=e_{j}P^{T}B^{*}(\lambda_{j})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(a_{i}\lambda_{j}),\,j=1,\ldots,N,$ (38)
where $e_{j}$, an $1\times N$ vector with the $j$th element equal to one and all the others equal to zero, and for $i,j=1,\ldots,N$:
[…]
Remark 12
It would be also of high importance to consider recursions that result in the following matrix functional equations:
$\tilde{Z}(s)=H(s)\sum_{i=1}^{N}Q^{(i)}\tilde{Z}(\zeta_{i}(s))+\tilde{V}(s).$
In our case $\zeta_{i}(s)=a_{i}s$, $i=1,\ldots,N$. It seems that for general $\zeta_{i}(s)$, the analysis can be handled similarly when we ensure that they are contractions on $\{s\in\mathbb{C}:Re(s)\geq 0\}$.
4.1 The case where $A_{n+1}$, $S_{n}$ are conditionally dependent based on the FGM copula
Contrary to the case considered above, we now assume that given $Y_{n}$, $Y_{n+1}$, the random variables $A_{n+1}$, $S_{n}$ are dependent based on the FGM copula. Under such an assumption, for $n\geq 0$, $x,y\geq 0$, $i,j=1,\ldots,N$:
$\begin{array}[]{l}P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i,A_{2},\ldots,A_{n},S_{1},\ldots,S_{n-1},Y_{1},\ldots,Y_{n-1})\vskip 5.69054pt\\ =P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i)=p_{i,j}F_{S,A|i,j}(y,x),\end{array}$ (39)
where, $F_{S,A|i,j}(y,x)$ is the bivariate distribution function of $(S_{n},A_{n+1})$ given $Y_{n}$, $Y_{n+1}$ with marginals $F_{S,i}(y):=B_{i}(y)$, $F_{A,j}(x)$ defined as $F_{S,A|i,j}(y,x)=C_{\Theta}^{FGM}(F_{S,i}(y),F_{A,j}(x))$ for $(y,x)\in\mathbb{R}^{+}\times\mathbb{R}^{+}$. The bivariate density of $(S,A)$ is given by
$\begin{array}[]{rl}f_{S,A|i,j}(y,x)=&c_{\Theta}^{FGM}(F_{S,i}(y),F_{A,j}(x))f_{S,i}(y)f_{A,j}(x)=f_{S,i}(y)f_{A,j}(x)+\theta_{i,j}g_{i}(y)(2\bar{F}_{A,j}(x)-1)f_{A,j}(x),\end{array}$
where $g_{i}(y):=f_{S,i}(y)(1-2F_{S,i}(y))$ with Laplace transform $g_{i}^{*}(s)=\int_{0}^{\infty}e^{-sy}g_{i}(y)dy$, $\bar{F}_{A,j}(x):=1-F_{A,j}(x)$, $f_{S,i}(y)$, $f_{A,j}(x)$ the densities of $S_{n}|Y_{n}=i$, $A_{n+1}|Y_{n+1}=j$, and $\theta_{i,j}\in[-1,1]$. In our case,
$\begin{array}[]{c}f_{S,A|i,j}(y,x)=f_{S,i}(y)\lambda_{j}e^{-\lambda_{j}x}+\theta_{i,j}g_{i}(y)\left[2\lambda_{j}e^{-2\lambda_{j}x}-\lambda_{j}e^{-\lambda_{j}x}\right].\end{array}$ (40)
Our aim is to obtain $Z_{j}(s;\Theta)=E(e^{-sW_{n}}1_{\{Y_{n}=j\}};\Theta)$, where $\Theta:=(\theta_{i,j})_{i,j=1,\ldots,N}$.
Theorem 13
The transforms $Z_{j}(s;\Theta)$, $j=1,\ldots,N$, satisfy the following equation:
$\begin{array}[]{rl}Z_{j}(s;\Theta)=&\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left(\beta_{i}^{*}(s)-\frac{\theta_{i,j}s}{2\lambda_{j}-s}g_{i}(s)\right)Z_{i}(a_{i}s;\Theta)\vskip 5.69054pt\\ &-s\sum_{i=1}^{N}p_{i,j}\left[\frac{\theta}{2\lambda_{j}-s}g_{i}(2\lambda_{j})Z_{i}(2a_{i}\lambda_{j};\Theta)+\frac{\beta_{i}^{*}(\lambda_{j})-\theta g_{i}(\lambda_{j})}{\lambda_{j}-s}Z_{i}(a_{i}\lambda_{j};\Theta)\right].\end{array}$ (41)
In matrix terms,
$\tilde{Z}(s;\Theta)=U(s;\Theta)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(a_{i}s;\Theta)+\tilde{V}(s;\Theta),$ (42)
where now,
$\begin{array}[]{rl}U(s;\Theta):=&L_{1}(s)P^{T}B^{*}(s)+(L_{1}(s)-L_{2}(s))(P^{T}\circ\Theta)G^{*}(s)\vskip 5.69054pt\\ =&H(s)+(L_{1}(s)-L_{2}(s))(P^{T}\circ\Theta)G^{*}(s),\vskip 5.69054pt\\ \tilde{V}(s;\Theta)=&(I-L_{1}(s))\tilde{v}^{(1)}+(I-L_{2}(s))\tilde{v}^{(2)},\end{array}$
with $P^{T}\circ\Theta$ denotes the $N\times N$ matrix with $(i,j)$ element equal to $\theta_{i,j}p_{i,j}$ (i.e., the operator $\circ$” denotes the Hadamard product), $G^{*}(s):=diag(g_{1}^{*}(s),\ldots,g_{N}^{*}(s))$, $L_{k}(s)=diag(\frac{k\lambda_{1}}{k\lambda_{1}-s},\ldots,\frac{k\lambda_{N}}{k\lambda_{N}-s})$, $\tilde{v}^{(k)}:=(v_{1}^{(k)},\ldots,v_{N}^{(k)})^{T}$, $k=1,2,$ where for $j=1,\ldots,N$,
$v_{j}^{(1)}:=\sum_{i=1}^{N}p_{i,j}\left(\beta_{i}^{*}(\lambda_{j})-\theta_{i,j}g_{i}^{*}(\lambda_{j})\right)Z_{i}(a_{i}\lambda_{j};\theta),\,\,\,v_{j}^{(2)}:=\sum_{i=1}^{N}p_{i,j}\theta_{i,j}g_{i}^{*}(2\lambda_{j})Z_{i}(2a_{i}\lambda_{j};\theta).$
The proof of Theorem 13 is similar to the one in Theorem 8 and further details are omitted. Following similar arguments as above, we have the following result.
Theorem 14
For $\tilde{V}(0;\Theta)=\mathbf{0}$,
[…]
where the functions $F_{i_{1},\ldots,i_{N}}(s;\Theta)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s;\Theta)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s;\Theta)U(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s);\Theta)\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(s;\Theta):=I$, and $F_{i_{1},\ldots,i_{N}}(s;\Theta)=O$ (i.e., the zero matrix) if one of the indices equals 1-1.
Now it remains to obtain the vectors $\tilde{v}^{(k)}$, $k=1,2,$ i.e., we need obtain $2N$ equations for the $2N$ unknowns $v_{j}^{(k)}$, $j=1,\ldots,N$, $k=1,2$. Setting $s=a_{i}\lambda_{j}$, and $s=2a_{i}\lambda_{j}$, $i,j=1,\ldots,N$ in (43) we obtain expressions for $Z_{i}(a_{i}\lambda_{j})$, $Z_{i}(a_{i}\lambda_{j})$.
Proposition 15
The vectors $\tilde{v}^{(k)}$, $k=1,2,$ are given as the unique solution of the following system of equations for $j=1,\ldots,N,$:
$\begin{array}[]{rl}v_{j}^{(1)}=&e_{j}[P^{T}B^{*}(\lambda_{j})+(P^{T}\circ\Theta)G^{*}(\lambda_{j})]\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(a_{i}\lambda_{j};\Theta),\vskip 5.69054pt\\ v_{j}^{(2)}=&e_{j}(P^{T}\circ\Theta)G^{*}(2\lambda_{j})]\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{Z}(2a_{i}\lambda_{j};\Theta),\end{array}$ (44)
where, for $m=1,2,$
[…]
Remark 16
Note that for $\Theta=O$, i.e., the independent copula with $\theta_{i,j}=0$, $i,j=1,\ldots,N,$ Theorem 14 is reduced to Theorem 10.
5 A modulated $D_{N}/G/1$ shot-noise queue
We now focus on the workload at arrival instants of a modulated $D_{N}/G/1$ shot-noise queue, that refers to a single server queue where the server’s speed is workload proportional, i.e., when the workload is $x$, the server’s speed equals $rx$; see [13] for a recent survey on shot-noise queueing systems, as well as [20, Section 6] that focused on the case where $N=1$. Our system operates as follows: Assume that the interarrival times $A_{1},A_{2},\ldots$ are such that $E(e^{-sA_{n}}1_{\{Y_{n-1}=i\}})=e^{-st_{i}}$, $i=1,\ldots,N$. There is a single server, and service requirements of successive customers $S_{1},S_{2},\ldots$ are i.i.d. random variables. We assume that just before the arrival of the $n$th customer, additional amount of work equal to $C_{n}$ is added. This can be explained as random noise caused by the arrival, and may be positive or negative. Let
$C_{n}=\left\{\begin{array}[]{ll}C_{n}^{+},&\text{ with probability }p,\\ -C_{n}^{-},&\text{ with probability }q=1-p.\end{array}\right$.
We further adopt the dynamics in (32), i.e.,, for $n\geq 0$, $x,y\geq 0$, $i,j=1,\ldots,N$:
$\begin{array}[]{l}P(C_{n+1}\leq x,S_{n}\leq y,Z_{n+1}=j|Z_{n}=i,(C_{r+1},S_{r},Z_{r}),r=0,1,\ldots,n-1)\vskip 5.69054pt\\ =P(C_{n+1}\leq x,S_{n}\leq y,Z_{n+1}=j|Z_{n}=i)=p_{i,j}F_{S,i}(y)G_{C,j}(x),\end{array}$
where $C^{+}|j$ has a general distribution with LST $c_{j}^{*}(s)$, and $C^{-}|j\sim\exp(\nu_{j})$, $j=1,\ldots,N$. Then, if $W_{n}$ is the workload before the $n$th arrival, we are dealing with a modulated stochastic recursion of the form $W_{n+1}=[e^{-rA_{n+1}}(W_{n}+S_{n})+C_{n+1}]^{+}$. Then, for $j=1,\ldots,N,$
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E\left(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}\right)=\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(e^{-sW_{n+1}}|Y_{n+1}=j,Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[pE\left(e^{-s(e^{-rA_{n+1}}(W_{n}+S_{n})+C^{+}_{n+1})}|Y_{n+1}=j,Y_{n}=i\right)\right.\vskip 5.69054pt\\ &\left.+qE\left(e^{-s[e^{-rA_{n+1}}(W_{n}+S_{n})-C^{-}_{n+1}]^{+}}|Y_{n+1}=j,Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[pc_{j}^{*}(-s)\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})\right.\vskip 5.69054pt\\ &\left.+qP(Y_{n}=i)E\left(\int_{y=0}^{e^{-rA_{n+1}(W_{n}+S_{n})}}e^{-s(e^{-rA_{n+1}}(W_{n}+S_{n})-y)}\nu_{j}e^{-\nu_{j}y}dy\right.\right.\vskip 5.69054pt\\ &\left.\left.+\int_{y=e^{-rA_{n+1}}(W_{n}+S_{n})}^{\infty}\nu_{j}e^{-\nu_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[pc_{j}^{*}(s)\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})\right.\vskip 5.69054pt\\ &\left.+qP(Y_{n}=i)E\left(\frac{\nu_{j}}{\nu_{j}-s}e^{-se^{-rA_{n+1}}(W_{n}+S_{n})}-\frac{s}{\nu_{j}-s}e^{\nu_{j}e^{-rA_{n+1}}(W_{n}+S_{n})}|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\left(pc_{j}^{*}(s)+q\frac{\nu_{j}}{\nu_{j}-s}\right)\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})-\frac{sq}{\nu_{j}-s}\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(\nu_{j}e^{-rt_{i}})Z_{i}^{n}(\nu_{j}e^{-rt_{i}}).\end{array}$
Letting $n\to\infty$, so that $Z_{j}^{n}(s)\to Z_{j}(s)$ we have
$Z_{j}(s)=\left(pc_{j}^{*}(s)+q\frac{\nu_{j}}{\nu_{j}-s}\right)\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(se^{-rt_{i}})Z_{i}(se^{-rt_{i}})-\frac{sq}{\nu_{j}-s}\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(\nu_{j}e^{-rt_{i}})Z_{i}(\nu_{j}e^{-rt_{i}}).$ (45)
In matrix notation, (45) is written as
$\tilde{Z}(s)=\tilde{C}(s)P^{T}\sum_{i=1}^{N}\tilde{P}^{(i)}B^{*}(se^{-rt_{i}})\tilde{Z}(se^{-rt_{i}})+\tilde{Q}(s),$ (46)
where $\tilde{P}^{(i)}$ as given in Theorem 10, $\tilde{C}(s):=pC(s)+q\widehat{L}(s)$, $C(s):=diag(c_{1}^{*}(s),\ldots,c_{N}^{*}(s))$, $\widehat{L}(s):=diag(\frac{\nu_{1}}{\nu_{1}-s},\ldots,\frac{\nu_{N}}{\nu_{N}-s})$, $\tilde{Q}(s):=q(I-\widehat{L}(s))\tilde{r}$, $\tilde{r}=(r_{1},\ldots,r_{N})$, with $r_{j}=\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(\nu_{j}e^{-rt_{i}})Z_{i}^{n}(\nu_{j}e^{-rt_{i}})$.
Note that (46) is more complicated with respect to (34), since the matrix $B^{*}(se^{-rt_{i}})$ is inside the summation, i.e., (46) is written as
$\tilde{Z}(s)=H(s)\sum_{i=1}^{N}\tilde{P}^{(i)}B^{*}(a_{i}s)\tilde{Z}(a_{i}s)+\tilde{Q}(s),$ (47)
where $a_{i}=e^{-rt_{i}}$, $i=1,\ldots,N$, $H(s):=\tilde{C}(s)P^{T}$. Moreover, the $n$th iterative of $\zeta_{i}(s):=a_{i}s$, i.e., $\zeta_{i}^{(n)}(s)=\zeta_{i}(\zeta_{i}(\ldots\zeta_{i}(s)\ldots))=se^{-rnt_{i}}\to 0$ as $n\to\infty$. Moreover, (47) is slightly different from (34), since the mappings $\zeta_{i}(s)$, $i=1,\ldots,N,$ are inside the summation. However, $H(e^{-rt_{i}m}s)\to P^{T}$ as $m\to\infty$, $i=1,\ldots,N$, $\tilde{Q}(0)=\mathbf{0}$. Thus, $\tilde{Z}(s)$ can be given similarly as in Theorem 10. Note that it remains to obtain the values of the vector $\tilde{r}$. This task can be accomplished using steps similar to those in Proposition 11, so further details are omitted.
Remark 17
Note that in this section we assumed that given $Y_{n}$, $Y_{n+1}$, $C_{n+1}$ and $S_{n}$ are conditionally independent. The analysis can be further applied in case we consider additional dependency, e.g., by assuming that $C_{n+1}$ and $S_{n}$ are dependent based on the FGM copula as in subsection 4.1, or have a (semi-)linear depend structure, e.g., $C_{n+1}=aS_{n}+J_{n}$, where $J_{n}$ independent random variable from $S_{n}$.
6 A modulated Markovian queue where service time depends on the waiting time
Consider the following modulated version of a variant of a M/M/1 queue that was investigated in [8, Section 5]; see also [20, subsection 6.2] for a modulated version. In particular, consider a variant of a M/M/1 queue, in which if the waiting time of the $n$th arriving customer equals $W_{n}$, then her service time equals $S_{n}-cW_{n}]^{+}$, where $c>0$. The dynamics in (32) are such that for $n\geq 0$, $x,y\geq 0$, $i,j=1,\ldots,N$:
$\begin{array}[]{l}P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i,A_{2},\ldots,A_{n},S_{1},\ldots,S_{n-1},Y_{1},\ldots,Y_{n-1})\vskip 5.69054pt\\ =P(A_{n+1}\leq x,S_{n}\leq y,Y_{n+1}=j|Y_{n}=i)=p_{i,j}(1-e^{-\mu_{i}y})(1-e^{-\lambda_{j}x}).\end{array}$
Contrary to the case in [20, subsection 6.2], we now assume that the service time also depends on the state of the background process.
Using similar arguments as above:
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E\left(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}\right)=\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(e^{-s[W_{n}+[S_{n}-cW_{n}]^{+}-A_{n+1}]^{+}}|Y_{n+1}=j,Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[E\left(\int_{0}^{W_{n}+[S_{n}-cW_{n}]^{+}}e^{-s(W_{n}+[S_{n}-cW_{n}]^{+}-y)}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right.\vskip 5.69054pt\\ &\left.+E\left(\int_{W_{n}+[S_{n}-cW_{n}]^{+}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}e^{-s(W_{n}+[S_{n}-cW_{n}]^{+})}-se^{-\lambda_{j}(W_{n}+[S_{n}-cW_{n}]^{+})}}{\lambda_{j}-s}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left(Z_{i}^{n}(s)-\frac{s}{\mu_{i}+s}Z_{i}^{n}(s+\mu_{i}c)\right)-\frac{s}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left(Z_{i}^{n}(\lambda_{j})-\frac{\lambda_{j}}{\mu_{i}+\lambda_{j}}Z_{i}^{n}(\lambda_{j}+\mu_{i}c)\right).\end{array}$
As $n\to\infty$, $Z^{n}_{i}(s)\to Z_{i}(s)$, we have,
$Z_{j}(s)=\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left[Z_{i}(s)-\frac{s}{\mu_{i}+s}Z_{i}(s+\mu_{i}c)\right]-\frac{s}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left[Z_{i}(\lambda_{j})-\frac{\lambda_{j}}{\mu_{i}+\lambda_{j}}Z_{i}(\lambda_{j}+\mu_{i}c)\right],$
or equivalently,
$\lambda_{j}\sum_{i=1}^{N}p_{i,j}Z_{i}(s)+(s-\lambda_{j})Z_{j}(s)-s\lambda_{j}\sum_{i=1}^{N}\frac{p_{i,j}}{\mu_{i}+s}Z_{i}(s+\mu_{i}c)=sv_{j},$ (48)
where $v_{j}=\sum_{i=1}^{N}p_{i,j}\left[Z_{i}(\lambda_{j})-\frac{\lambda_{j}}{\mu_{i}+\lambda_{j}}Z_{i}(\lambda_{j}+\mu_{i}c)\right]$, $j=1,\ldots,N$.
For $s=0$, (48) yields $Z_{j}(0)=\sum_{i=1}^{N}p_{i,j}Z_{i}(0)$, thus, $\tilde{Z}(0)=\tilde{\pi}$. In matrix terms, (48) is rewritten as:
$D(s)\tilde{Z}(s)=s\tilde{v}+\Lambda P^{T}\sum_{i=1}^{N}H^{(i)}(s)\tilde{Z}(s+\mu_{i}c),$ (49)
where $D(s)=sI-\Lambda(I-P^{T})$, $\tilde{v}=(v_{1},\ldots,v_{N})^{T}$, $H^{(i)}(s)=(I-M(s))\tilde{P}^{(i)}$, $M(s)=diag(\frac{\mu_{1}}{\mu_{1}+s},\ldots,\frac{\mu_{N}}{\mu_{N}+s})$. Note that $H^{(i)}(s)$ is an $N\times N$ matrix with the $(i,i)$ element equal to $\frac{1}{\mu_{i}+s}$, $i=1,2,\ldots,N$, and all other elements equal to zero. Note that $v_{j}=P(W=0;1_{\{Y_{n+1}=j\}})$.
Lemma 18
The matrix $\Lambda(I-P^{T})$ has exactly $N$ eigenvalues $\gamma_{i}$, $i=1,\ldots,N$, with $\gamma_{1}=0$, and $Re(\gamma_{i})>0$, $i=2,\ldots,N$.
Proof. Clearly, $s:=\gamma_{1}=0$ is a root of $det(D(s)=0)$, since $P$ is a stochastic matrix. By applying Gersgorin’s circle theorem [25, Th. 1, Section 10.6], every eigenvalue of $\Lambda(I-P^{T})$ lies in at least one of the disks
$\{s:|s-\lambda_{i}(1-p_{i,i})|\leq\sum_{k\neq i}|\lambda_{i}p_{k,i}|=\lambda_{i}\sum_{k\neq i}p_{k,i}\}.$
Therefore, for each $i$, the real part of $\gamma_{i}$ is positive.
Let,
$\zeta_{s}:=\{s:Re(s)\geq 0,Det(D(s))\neq 0\}.$
Then, for $s\in\zeta_{s}$,
$\tilde{Z}(s)=A(s)\tilde{v}+G(s)\sum_{i=1}^{N}H^{(i)}(s)\tilde{Z}(s+\mu_{i}c),$ (50)
where $A(s):=sD^{-1}(s)$, $G(s):=D^{-1}(s)\Lambda P^{T}$. Iterating (50) and having in mind that $\tilde{Z}(s)\to\mathbf{0}$ as $s\to\infty$,
$\tilde{Z}(s)=\sum_{k=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=k}L_{i_{1},\ldots,i_{N}}(s)A(\zeta_{i_{1},\ldots,i_{N}}(s))\tilde{v}$, (51)
where the functions $L_{i_{1},\ldots,i_{N}}(s)$ are derived recursively by,
$L_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}L_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)G(\zeta_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))H^{(k)}(\zeta_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)),$
and
$\zeta_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)=\zeta_{1}^{i_{1}}(\zeta_{2}^{i_{3}}(\ldots(\zeta_{N}^{i_{N}}(s))\ldots),$
and $\zeta_{k}^{i_{k}}(s)$ is the $k$th iterate of $\zeta_{k}(s)=s+\mu_{k}c$, i.e., $\zeta_{k}^{i_{k}}(s)=s+i_{k}\mu_{k}c$, $k=1,\ldots,N$.
Note that $D(s)$ is singular at the eigenvalues of $\Lambda(I-P^{T})$, i.e., $det(D(s))=0$ at $s=\gamma_{l}$, $l=1,\ldots,N$. However, $\tilde{Z}(s)$ is analytic in the half-plane $Re(s)\geq 0$, and thus, the vector $\tilde{v}$ will be derived so that the right-hand side of (51) is finite at $s=\gamma_{l}$, $l=1,\ldots,N$. Divide (49) with $s$ and denote by $\tilde{y}_{l}$, the left (row) eigenvector of $\Lambda(I-P^{T})$, associated with the eigenvalue $\gamma_{l}$, $l=1,\ldots,N$ ($y_{1}=\mathbf{1}$ is the row eigenvector with all elements equal to 1, corresponding to the eigenvalue $\gamma_{1}=0$). Then, (49) is written as
$\tilde{y}_{l}(1-\frac{\gamma_{l}}{s})\tilde{Z}(s)=\tilde{v}+\tilde{y}_{l}\Lambda P^{T}\sum_{i=1}^{N}\tilde{T}^{(i)}(s)\tilde{Z}(s+\mu_{i}c),\,l=1,\ldots,N,\,Re(s)\geq 0,$ (52)
where $\tilde{T}^{(i)}(s):=s^{-1}H^{(i)}(s)=diag(\frac{1}{\mu_{1}+s},\ldots,\frac{1}{\mu_{N}+s})\tilde{P}^{(i)}$, $i=1,\ldots,N$.
Letting $s=\gamma_{l}$, $l=1,\ldots,N$, and using (51), we obtain $N$ equations for the derivation of the $N$ elements of $\tilde{v}$:
$\tilde{y}_{l}\tilde{v}=1_{\{l=1\}}+\frac{1}{\mu+\gamma_{i}}\tilde{y}_{i}\Lambda P^{T}\sum_{i=1}^{N}\tilde{T}^{(i)}(\gamma_{l})\sum_{k=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=k}L_{i_{1},\ldots,i_{N}}(\gamma_{l}+\mu_{i}c)A(\zeta_{i_{1},\ldots,i_{N}}(\gamma_{l}+\mu_{i}c))\tilde{v}$. (53)
7 A modulated ASIP tandem network with consumption
In this section, we consider the Markov-modulated analogue of the model analyzed in [9]. The authors in [9] considered a non-modulated ASIP (asymmetric inclusion process) tandem queue, in which the first queue receives a fluid input according to a Lévy subordinator process. Each queue has a gate that opens after independent, exponentially distributed periods for an infinitesimal amount of time, allowing the queue content to move to the next queue. In addition, again at independent exponentially distributed instants, a fixed fraction of a queue content is removed from the system.
Consider a queueing model consisting of two stations, say $Q_{1},Q_{2}$ in series. We assume that only $Q_{1}$ receives an external input, which is a Lévy subordinator process $X=\{X(t);t\geq 0\}$. Given the state of the background process $J(t)$, that takes values in $E=\{1,2,\ldots,N\}$, the Laplace exponent of $\{X(t);t\geq 0\}$ is $\phi_{i}(.)$, with $\phi_{i}(0)=0$, i.e., That is, $E(e^{-sX(t)}1_{\{J(t)=i\}})=e^{-\phi_{i}(s)t}$ for $s,t\geq 0$. The process $\{J(t);t\geq 0\}$ jumps to any state $j\in E$ with probability $p_{i,j}$, given that $J(t)=i$. In state $i\in E$, $J(t)$ remains for an exponentially distributed time interval with rate $q_{i}$. Given that $J(t)=i$, $Q_{k}$ has a gate that is closed except for infinitesimally short gate openings that occur at independent $exp(\mu_{k,i})$ intervals, $k=1,2$. At a gate opening of $Q_{1}$, its content instantaneously moves to $Q_{2}$; at a gate opening of $Q_{2}$, its content leaves the system. At independent $exp(\tau_{k,i})$ intervals, a fraction $a_{k,i}$ of the content of station $Q_{k}$ is instantaneously removed from the system, $k=1,2$. Let $(Z_{1}(t),Z_{2}(t))$ be the buffer contents of $(Q_{1},Q_{2})$ at time $t\geq 0$, with $Z_{1}(0)=0=Z_{2}(0)$, and with LST $f_{i}(t,s_{1},s_{2})=E(e^{-s_{1}Z_{1}(t)-s_{2}Z_{2}(t)}1_{\{J(t)=i\}})$, $i\in E$. Then, for $h\to 0$ we have,
$\begin{array}[]{rl}f_{i}(t+h,s_{1},s_{2})=&[1-(\mu_{1,i}+\mu_{2,i}+\tau_{1,i}+\tau_{2,i}+q_{i})h]f_{i}(t,s_{1},s_{2})+\mu_{1,i}hf_{i}(t,s_{2},s_{2})+\mu_{2,i}hf_{i}(t,s_{1},0)\vskip 5.69054pt\\ &+\tau_{1,i}hf_{i}(t,(1-a_{1,i})s_{1},s_{2})+\tau_{2,i}hf_{i}(t,s_{1},(1-a_{2,i})s_{2})+q_{i}h\sum_{j=1}^{N}p_{i,j}f_{j}(t,s_{1},s_{2})+o(h).\end{array}$
For the steady-state case, with $f_{i}(s_{1},s_{2})$ the LST of the steady-state joint buffer content distribution, we have
$\begin{array}[]{l}(\mu_{1,i}+\mu_{2,i}+\tau_{1,i}+\tau_{2,i}+q_{i})f_{i}(s_{1},s_{2})=\mu_{1,i}f_{i}(s_{2},s_{2})+\mu_{2,i}f_{i}(s_{1},0)\vskip 5.69054pt\\ +\tau_{1,i}f_{i}((1-a_{1,i})s_{1},s_{2})+\tau_{2,i}f_{i}(s_{1},(1-a_{2,i})s_{2})+q_{i}\sum_{j=1}^{N}p_{i,j}f_{j}(s_{1},s_{2}).\end{array}$ (54)
Let $\tilde{f}(s_{1},s_{2})=(f_{1}(s_{1},s_{2}),\ldots,f_{N}(s_{1},s_{2}))^{T}$, $M_{k}=diag(\mu_{k,1},\ldots,\mu_{k,N})$, $T_{k}=diag(\tau_{k,1},\ldots,\tau_{k,N})$, $k=1,2,$ $Q=(q_{i,j})_{i,j\in E}$, with $q_{i,j}:=q_{i}p_{i,j}$, $\in E, iji\neq j$, $q_{i,i}:=q_{i}(p_{i,i}-1)$, $i\in E$, and $\tilde{\phi}(s_{1})=diag(\phi_{1}(s_{1}),\ldots,\phi_{N}(s_{1}))$, then, (54) can be rewritten in matrix form as
$\begin{array}[]{c}(M_{1}+M_{2}+T_{1}+T_{2}+\tilde{\phi}(s_{1})-Q)\tilde{f}(s_{1},s_{2})=M_{1}\tilde{f}(s_{2},s_{2})+M_{2}\tilde{f}(s_{1},0)\vskip 5.69054pt\\ +T_{1}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s_{1},s_{2})+T_{2}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{1},(1-a_{2,i})s_{2}).\end{array}$ (55)
7.1 The buffer content of $Q_{1}$
Focusing on the marginal content of $Q_{1}$, we have that the LST $\tilde{f}(s):=\tilde{f}(s,0)$, and $\tilde{f}(0)=\tilde{\pi}^{T}$, the stationary vector of the background Markov process $\{J(t);t\geq 0\}$. Therefore, (55) is reduced to
$(M_{1}+T_{1}+\tilde{\phi}(s)-Q)\tilde{f}(s)=M_{1}\tilde{\pi}^{T}+T_{1}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s),\,Re(s)\geq 0.$
Then, we have the following result.
Lemma 19
The equation
$det(M_{1}+T_{1}+\tilde{\phi}(s)-Q)=0,$ (56)
has exactly $N$ roots, say $y_{1},\ldots,y_{N}$, such that $Re(y_{i})<0$, $i=1,\ldots,N$.
Proof. This is easily proven by writing first $Q=\bar{Q}-Q_{d}$, where $\bar{Q}=(q_{i}p_{i,j})_{i,j\in E}$, $Q_{d}:=diag(q_{1},\ldots,q_{N})$. Then, $G(s):=M_{1}+T_{1}+\tilde{\phi}(s)-Q=F(s)-\bar{Q}$, where $F(s):=diag(f_{1}(s),\ldots,f_{N}(s))$, $f_{i}(s):=\mu_{1,i}+\tau_{1,i}+q_{i}+\phi_{i}(s)$, $i\in E$. Then,
$|\mu_{1,i}+\tau_{1,i}+q_{i}+\phi_{i}(s)|>\sum_{j=1}^{N}|q_{i}p_{i,j}|=q_{i}.$
By using [17, Theorem 11.3], since $det(F(s))=0$ has exactly $N$ roots with a negative real part, is implied that $det(F(s)-\bar{Q})=det(M_{1}+T_{1}+\tilde{\phi}(s)-Q)=0$ has exactly $N$ roots with a negative real part.
Therefore,
$\tilde{f}(s)=\tilde{M}(s)+\tilde{T}(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s),\,Re(s)\geq 0,$ (57)
where, $\tilde{M}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}M_{1}\tilde{\pi}^{T}$, $\tilde{T}(s):=(M_{1}+T_{1}+\tilde{\phi}(s)-Q)^{-1}T_{1}$. Note that (57) has exactly the same form as (34). Iteration of (57) yields
$\tilde{f}(s)=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}F_{i_{1},\ldots,i_{N}}(s)\tilde{M}(a_{i_{1},\ldots,i_{N}}(s)),$ (58)
where now, $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=(1-a_{1,i})s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(s)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)\tilde{T}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(s):=I$, and $F_{i_{1},\ldots,i_{N}}(s)=O$ (i.e., the zero matrix) if one of the indices equals 1-1. Note that (58) is derived by having in mind that as $n\to\infty$, $a_{i_{1},\ldots,i_{N}}(s)\to 0$, and $\tilde{T}(a_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}T_{1}$, and $\tilde{M}(a_{i_{1},\ldots,i_{N}}(s))\to(M_{1}+T_{1}-Q)^{-1}M_{1}\tilde{\pi}^{T}$. Now, by iterating (57) $n$ times yields,
$\tilde{f}(s)=\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(s)\tilde{M}(a_{i_{1},\ldots,i_{N}}(s))+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(s)\tilde{f}(a_{i_{1},\ldots,i_{N}}(s)),$ (59)
Note that $F_{i_{1},\ldots,i_{N}}(s)$ is written as the sum of the product of matrices, where each has elements that are strictly smaller than one. This implies that the second term on the right-hand side of (59) is a convergent matrix as $n\to\infty$. Thus, we come up with (58).
Remark 20
For $a_{1,i}=0$, $i=1,\ldots,N$, (58) becomes
$\tilde{f}(s)=(I-\tilde{T}(s))^{-1}\tilde{M}(s),$ (60)
and when $a_{1,i}=1$, $i=1,\ldots,N$, we have $\tilde{f}(s)=\tilde{M}(s)$.
Let us now consider (55) $a_{2,i}=0$, $a_{1,i}\in(0,1)$, $i\in E$. Then, (55) is reduced to
$(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)\tilde{f}(s_{1},s_{2})=M_{1}\tilde{f}(s_{2},s_{2})+M_{2}\tilde{f}(s_{1},0)+T_{1}\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s_{1},s_{2}),$
or equivalently to
$\begin{array}[]{c}\tilde{f}(s_{1},s_{2})=N_{1}(s_{1})\tilde{f}(s_{2},s_{2})+N_{2}(s_{1})\tilde{f}(s_{1},0)+T_{1,1}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s_{1},s_{2}),\end{array}$ (61)
where
$\begin{array}[]{rl}N_{i}(s_{1}):=&(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)^{-1}M_{i},\,i=1,2,\\ T_{1,1}(s_{1}):=&(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)^{-1}T_{1}.\end{array}$
Remark 21
One can easily show (as in Lemma 19) that $det(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)=0$ has NN roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{1}+\tilde{\phi}(s_{1})-Q)^{-1}$ exits.
Set $L(s_{1},s_{2}):=N_{1}(s_{1})\tilde{f}(s_{2},s_{2})+N_{2}(s_{1})\tilde{f}(s_{1},0)$ and (61) is rewritten for $Re(s_{1})\geq 0,Re(s_{2})\geq 0$ as
$\tilde{f}(s_{1},s_{2})=L(s_{1},s_{2})+T_{1,1}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}((1-a_{1,i})s_{1},s_{2}).$ (62)
Note that for fixed $s_{2}$ such that $Re(s_{2})\geq 0$, (62) has the same form as (61). Note that as $n\to\infty$, $a_{i_{1},\ldots,i_{N}}(s_{1})\to 0$, $N_{i}(a_{i_{1},\ldots,i_{N}}(s_{1}))\to(M_{1}+M_{2}+T_{1}-Q)^{-1}M_{i}$, i.e., the elements of the limiting matrix are all bounded by a number that is smaller than 1. Therefore, under similar arguments as in (58),
$\tilde{f}(s_{1},s_{2})=\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2}),$ (63)
where now, $a_{i_{1},\ldots,i_{N}}(s_{1})=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s_{1}))\ldots))$ and $a_{i}^{n}(s_{1})$ denotes the $n$th iterate of $a_{i}(s_{1})=(1-a_{1,i})s_{1}$, with $a_{0,\ldots,0}(s_{1})=s_{1}$, and the functions $\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})$ are recursively defined by
$\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})=\sum_{k=1}^{N}\tilde{F}_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1},s_{2})T_{1,1}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s_{1}))\tilde{P}^{(k)},$
with $\tilde{F}_{0,\ldots,0}(s_{1},s_{2}):=I$, and $\tilde{F}_{0,\ldots,0}(s_{1}),s_{2})=O$ (i.e., the zero matrix) if one of the indices equals 1-1. Note also that the vector $\tilde{f}(s_{1},0)$ in the expression of $L(s_{1},s_{2})$ has already been derived in (58). Thus, $\tilde{f}(a_{i_{1},\ldots,i_{N}}(s_{1}),0)$ that is needed in $L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ in (63) is already known. We further need to derive an expression for $\tilde{f}(s_{2},s_{2})$ so that $L(a_{i_{1},\ldots,i_{N}}(s_{1}),s_{2})$ to be fully determined. Set $s_{1}=s_{2}$ in (63), and obtain after simple calculations:
$\begin{array}[]{l}[I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{1}(a_{i_{1},\ldots,i_{N}}(s_{1}))]\tilde{f}(s_{2},s_{2})\vskip 5.69054pt\\ =\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{2}(a_{i_{1},\ldots,i_{N}}(s_{1}))\tilde{f}(a_{i_{1},\ldots,i_{N}}(s_{1}),0).\end{array}$
Provided that the $I-\sum_{j=0}^{\infty}\sum_{i_{1}+i_{2}+\ldots+i_{N}=j}\tilde{F}_{i_{1},\ldots,i_{N}}(s_{1},s_{2})N_{1}(a_{i_{1},\ldots,i_{N}}(s_{1}))$ is invertible, we can have an expression for $\tilde{f}(s_{2},s_{2})$. Therefore, we have obtained $\tilde{f}(s_{1},s_{2})$ through (63).
7.2 The case $a_{1,i}=0$, $i=1,\ldots,N$
In this case, first note that $\tilde{f}(s_{1},0)$ is given by (63). Then, (55) is written as
$\tilde{f}(s_{1},s_{2})=\widehat{M}_{1}(s_{1})\tilde{f}(s_{2},s_{2})+\widehat{M}_{2}(s_{1})\tilde{f}(s_{1},0)+\widehat{T}_{2}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{1},(1-a_{2,i})s_{2}),$ (64)
where $\widehat{M}_{i}(s_{1}):=(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)^{-1}M_{i}$, $i=1,2$ and $\widehat{T}_{2}(s_{1}):=(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)^{-1}T_{2}$.
Remark 22
One can easily show (as in Lemma 19) that $det(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)=0$ has NN roots with negative real parts, thus, for $Re(s)\geq 0$, $(M_{1}+M_{2}+T_{2}+\tilde{\phi}(s_{1})-Q)^{-1}$ exits.
Substituting $s_{1}=s_{2}$ we obtain
$(I-\widehat{M}_{1}(s_{2}))\tilde{f}(s_{2},s_{2})=\widehat{M}_{2}(s_{1})\tilde{f}(s_{2},0)+\widehat{T}_{2}(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2}),$
or equivalently,
$\tilde{f}(s_{2},s_{2})=\widehat{F}(s_{2})\tilde{f}(s_{2},0)+\widehat{G}(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2}),$ (65)
where $\widehat{F}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{M}(s_{2})$, $\widehat{G}(s_{2}):=(I-\widehat{M}_{1}(s_{2}))^{-1}\widehat{T}(s_{2})$. Substituting (63), we get an expression for $\tilde{f}(s_{2},s_{2})$. Substituting back in (64) we come up with the following relation
$\begin{array}[]{rl}\tilde{f}(s_{1},s_{2})=&\widehat{M}_{1}(s_{1})\widehat{F}(s_{2})\tilde{f}(s_{2},0)+\widehat{M}_{1}(s_{1})G(s_{2})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{2},(1-a_{2,i})s_{2})\vskip 5.69054pt\\ &+\widehat{M}_{2}(s_{1})\tilde{f}(s_{1},0)+\widehat{T}_{2}(s_{1})\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s_{1},(1-a_{2,i})s_{2}),\end{array}$ (66)
where thanks to (63), $\tilde{f}(s,0)=(I-\tilde{T}(s))^{-1}\tilde{M}(s)$. Replacing $s_{1}$ with $s$, $s_{2}$ with $t$, and writing $b_{i}:=1-a_{2,i}$, $i=1,\ldots,N$, (66) is rewritten as
$\tilde{f}(s,t)=K(s,t)+R(s,t)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(t,b_{i}t)+\widehat{T}_{2}(s)\sum_{i=1}^{N}\tilde{P}^{(i)}\tilde{f}(s,b_{i}t),$ (67)
where,
$\begin{array}[]{rl}K(s,t):=&\widehat{M}_{1}(s)\widehat{F}(t)(I-\tilde{T}(t))^{-1}\tilde{M}(t)+\widehat{M}_{2}(s)(I-\tilde{T}(s))^{-1}\tilde{M}(s),\\ R(s,t):=&\widehat{M}_{1}(s)\widehat{G}(t).\end{array}$
Iterating (67) n1n-1 times results in
$\begin{array}[]{rl}\tilde{f}_{n}(s,t)=&K_{n}(s,t)+\sum_{k=1}^{n}\sum_{i_{1}+\ldots+i_{N}=k-1}\sum_{j_{1}+\ldots+j_{N}=n,,j_{l}\geq i_{l}}G_{\tilde{i},\tilde{j}}(s,t)\tilde{f}(b_{i_{1},\ldots,i_{N}}(t),b_{j_{1},\ldots,j_{N}}(t))\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n}F_{i_{1},\ldots,i_{N}}(s)\tilde{f}(s,b_{i_{1},\ldots,i_{N}}(t)),\end{array}$ (68)
where, $j_{l}\geq i_{l}$, $l=1,\ldots,N$, and $\tilde{i}:=(i_{1},\ldots,i_{N})$
$\begin{array}[]{rl}K_{n}(s,t)=&K_{n-1}(s,t)+\sum_{k=1}^{n-1}\sum_{i_{1}+\ldots+i_{N}=k-1}\sum_{j_{1}+\ldots+j_{N}=n-1,j_{l}\geq i_{l}}G_{\tilde{i},\tilde{j}}(s,t)K(b_{i_{1},\ldots,i_{N}}(t),b_{j_{1},\ldots,j_{N}}(t))\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n-1}F_{i_{1},\ldots,i_{N}}(s)K(s,b_{i_{1},\ldots,i_{N}}(t)),\,n=2,3,\ldots,\vskip 5.69054pt\\ F_{i_{1},\ldots,i_{N}}(s)=&\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)T_{1}(s)\tilde{P}^{(k)},\end{array}$
and for $j_{1}+\ldots+j_{N}=n$,
$\begin{array}[]{rl}G_{\tilde{i},\tilde{j}}(s,t)=&\left[\sum_{i_{1}+\ldots+i_{N}\leq n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)G_{\tilde{0},\tilde{1}_{k}}(b_{\tilde{i}}(t),b_{\tilde{j}-1_{k}}(t),t))+F_{\tilde{i}}(s)G_{\tilde{0},\tilde{j}-\tilde{i}}(s,b_{\tilde{i}}(t))\right]1_{\{i_{1}+\ldots+i_{N}=n-1\}}\vskip 5.69054pt\\ &+\sum_{j_{1}+\ldots+j_{N}=n-1,j_{l}\geq i_{l}}\sum_{k=1}^{N}G_{\tilde{i},\tilde{j}-\tilde{1}_{k}}(s,t)F_{\tilde{1}_{k}}(b_{\tilde{i}}(t))1_{\{i_{1}+\ldots+i_{N}<n-1\}}.\end{array}$
In case $\mu_{2,i}>0$, one can approximate $\tilde{f}(s,t)$ arbitrarily closely by $K_{n}(s,t)$, by taking $n$ sufficiently large. This is because the second and third terms in the righthand side of (68) become arbitrarily small as $n\to\infty$.
8 An integer vector-valued reflected autoregressive process
Consider an integer-valued stochastic process recursion that is described by
$X_{n+1}=[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{+}.$ (69)
Such a recursion describes the number of waiting customers in a generalized M/G/1 queue with impatient customers, just after the beginning of the $n$th service. $X_{n}$ describes that number and $B_{n}$ is the number of customers arriving during the service time of the $n$th customer. The service times are governed by a Markov process $Z_{n}$, $n=0,1,\ldots$ that takes values in $E=\{1,2,...,N\}$. $U_{k,n}$ are i.i.d. Bernoulli distributed random variables with $P(U_{k,n}=1|Z_{n}=i)=\xi_{n}^{(i)}$, $P(U_{k,n}=0|Z_{n}=i)=1-\xi_{n}^{(i)}$. Moreover, we assume that the $\xi_{n}^{(i)}$ are themselves random variables, independent and identically distributed with $P(\xi_{n}^{(i)}=a_{i,j}|Z_{n}=i)=q_{i,j}$ , i,j=1,,Ni,j=1,...,N, with $a_{i,j}\in(0,1)$. Moreover, set $Q=(q_{i,j})_{i,j=1,\ldots,N}$.
Denote by for i,j=1,2,,Ni,j=1,2,...,N, $|z|\leq 1$,
$\begin{array}[]{rl}B_{i,j}(z):=&E(z^{B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i),\end{array}$
and let $B(z)=(B_{i,j}(z))_{i,j=1,\ldots,N}$. Then,
$\begin{array}[]{rl}E(z^{X_{n+1}}1_{\{Z_{n+1}=j\}})=&E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{+}}1_{\{Z_{n+1}=j\}})\vskip 5.69054pt\\ =&E(\left(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}+1-z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}\right)1_{\{Z_{n+1}=j\}})\vskip 5.69054pt\\ =&E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}1_{\{Z_{n+1}=j\}})+P(Z_{n+1}=j)-E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}1_{\{Z_{n+1}=j\}}).\end{array}$ (70)
Now,
$\begin{array}[]{rl}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1}1_{\{Z_{n+1}=j\}})=&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i)P(Z_{n}=i)\vskip 5.69054pt\\ =&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)E(z^{B_{n}}1_{\{Z_{n+1}=j\}}|Z_{n}=i)P(Z_{n}=i)\vskip 5.69054pt\\ =&\frac{1}{z}\sum_{i=1}^{N}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)B_{i,j}(z)P(Z_{n}=i).\end{array}$ (71)
Then, tedious but standard calculations yield,
$\begin{array}[]{rl}E(z^{\sum_{k=1}^{X_{n}}U_{k,n}}|Z_{n}=i)P(Z_{n}=i)=&\sum_{l=1}^{N}q_{i,l}E((a_{i,j}(z))^{X_{n}}1_{\{Z_{n}=i\}}),\end{array}$ (72)
where $a_{i,j}(z):=\bar{a}_{i,j}+a_{i,j}z$, $\bar{a}_{i,j}:=1-a_{i,j}$, $i,j=1,\ldots,N$. Note that $a_{i,j}(z)$, $i,j=1,\ldots,N$, are commutative contraction mappings on the closed unit disk. Moreover,
$\begin{array}[]{rl}E(z^{[\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1]^{-}}1_{\{Z_{n+1}=j\}})=&E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1\geq 0\}})+\frac{1}{z}E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1=-1\}})\vskip 5.69054pt\\ =&P(Z_{n+1}=j)-(1-\frac{1}{z})E(1_{\{Z_{n+1}=j\}}1_{\{\sum_{k=1}^{X_{n}}U_{k,n}+B_{n}-1=-1\}}).\end{array}$ (73)
Denoting $f_{j}(z)=\lim_{n\to\infty}E(z^{X_{n}}1_{\{Z_{n}=j\}})$, $\tilde{F}(z):=(f_{1}(z),\ldots,f_{N}(z))^{T}$, we have the following result.
Theorem 23
The generating functions $f_{j}(z)$, $j=1,\ldots,N$, satisfy the following system
$\begin{array}[]{rl}f_{j}(z)=&\frac{1}{z}\sum_{i=1}^{N}B_{i,j}(z)\sum_{l=1}^{N}q_{i,l}f_{i}(a_{i,l}(z))+(1-\frac{1}{z})q_{-1,j},\end{array}$ (74)
or equivalently, in matrix notation
$\tilde{F}(z)=\frac{1}{z}B^{T}(z)\sum_{i=1}^{N}Q^{(i)}\sum_{l=1}^{N}P_{i}^{(l)}\tilde{F}(a_{i,l}(z))+\tilde{K}(z),$ (75)
where $\tilde{K}(z)=(1-\frac{1}{z})\tilde{q}_{-1}$, $\tilde{q}_{-1}:=(q_{-1,1},\ldots,q_{-1,N})^{T}$, and $q_{-1,j}=\sum_{i=1}^{N}B_{i,j}(0)\sum_{l=1}^{N}q_{i,l}f_{i}(\bar{a}_{i,l})$, $j=1,\ldots,N$. Moreover, $Q^{(i)}$ is a $N\times N$ matrix with rows equal to zero except row $i$ that coincides with row $i$ of matrix $Q$, $P_{l}^{(i)}$ is a $N\times N$ matrix with all entities equal to zero except the $(i,l)$ entity which is equal to one. Note that $\sum_{i=1}^{N}Q^{(i)}\sum_{l=1}^{N}P_{i}^{(l)}=I$.
Proof. Substituting (71)-(73) in (70), and letting $n\to\infty$ we obtain after tedious calculations in (74). Now multiplying (74) with $z$ and then letting $z=0$, we obtain the expression for $q_{-1,j}$. In matrix notation, (74) is written as (75).
Setting $G(z):=\frac{1}{z}B^{T}(z)$ and $T_{i,j}=Q^{(i)}P_{i}^{(j)}$, $i,j=1,\ldots,N$, (75) is rewritten as
$\tilde{F}(z)=G(z)\sum_{i=1}^{N}\sum_{j=1}^{N}T_{i,j}\tilde{F}(a_{i,j}(z))+\tilde{K}(z),$ (76)
Note that the fixed point of the iterates $a_{i,j}(z)=1-a_{i,j}+a_{i,j}z$ is $z=1$, and we have that $\tilde{K}(1)=\textbf{0}$. Thus, $\tilde{F}(z)$ follows from a modification of Theorem 10. In particular, let
$a^{i(1,1),i(1,2),\ldots,(N,N)}(z):=a_{1,1}^{i(1,1)}(a_{1,2}^{i(1,2)}(\ldots a_{N,N}^{i(N,N)}(z)\ldots)),$
and $a_{m,l}^{n}(z)$ is defined as the $n$th iterate of $a_{m,l}(z)$ with $a^{0,0,\ldots,0}(z)=z$. Iterating $n$ times (76) yields
$\begin{array}[]{rl}\tilde{F}(z)=&\sum_{\sum_{l,m=1}^{N}i(l,m)=n+1}L_{i(1,1),i(1,2),\ldots,(N,N)}(z)\tilde{F}(a^{i(1,1),i(1,2),\ldots,i(N,N)}(z))\vskip 5.69054pt\\ &+\sum_{k=0}^{n}\sum_{\sum_{l,m=1}^{N}i(l,m)=k}L_{i(1,1),i(1,2),\ldots,i(N,N)}(z)\tilde{K}(a^{i(1,1),i(1,2),\ldots,(N,N)}(z)),\end{array}$ (77)
where the matrix functions $L_{i(1,1),i(1,2),\ldots,i(N,N)}(z)$ are derived recursively by
$L_{i(1,1),i(1,2),\ldots,i(N,N)}(z)=\sum_{u=1}^{N}\sum_{v=1}^{N}L_{i(1,1),\ldots,i(u,v)-1,\ldots,i(N,N)}(z)G(a^{i(1,1),\ldots,i(u,v)-1,\ldots,(N,N)})T_{v,l},$ (78)
with $L_{0,0,\ldots,0}(z)=I$. Using similar arguments as in Theorem 10 we have
$\begin{array}[]{rl}\tilde{F}(z)=&\lim_{n\to\infty}\sum_{\sum_{l,m=1}^{N}i(l,m)=n+1}L_{i(1,1),i(1,2),\ldots,(N,N)}(z)\tilde{\pi}\vskip 5.69054pt\\ &+\sum_{k=0}^{\infty}\sum_{\sum_{l,m=1}^{N}i(l,m)=k}L_{i(1,1),i(1,2),\ldots,i(N,N)}(z)\tilde{K}(a^{i(1,1),i(1,2),\ldots,(N,N)}(z)),\end{array}$ (79)
We still need to derive $\tilde{q}_{-1}$. This can be done by substituting $z=\bar{a}_{i,j}$ in (79) and substituting the $j$th component of the derived $\tilde{F}(\bar{a}_{i,j})$ in the expression for $q_{-1,j}$ given in Theorem 23.
9 On a reflected VAR(1) process
We now consider the a reflected vector autoregressive process (VAR(1)) (see e.g., [23, 21], [14, Chapter 4]) that is described by the following vector-valued recursion:
$\tilde{Z}_{n+1}=[A\tilde{Z}_{n}+\tilde{S}_{n}-\tilde{A}_{n+1}]^{+},$ (80)
where $A$ is an $N\times N$ matrix with elements $a_{i,j}$, $i,j=1,\ldots,N$, such that $a_{i,j}\in[0,1)$, and $\tilde{Z}_{n}:=(Z_{1,n},\ldots,Z_{N,n})^{T}$. Let also $\tilde{S}_{n}:=(S_{1,n},\ldots,S_{N,n})^{T}$, $\tilde{A}_{n}:=(A_{1,n},\ldots,A_{N,n})^{T}$, and assume that $S_{j,n}$, $A_{j,n}$ are series of i.i.d. random variables, independent of anything else. From (80) is readily seen that
$Z_{j,n+1}=[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{+},\,j=1,\ldots,N.$ (81)
9.1 The vector of the marginal Laplace-Stieltjes transforms
Let $F^{n}_{j}(s):=E(e^{-sF_{j,n}})$ be the LST of the distribution of a random variable $F_{j,n}$, and $\tilde{Z}^{n}(s)=(Z_{1}^{n}(s),\ldots,Z_{N}^{n}(s))^{T}$. Then,
$\begin{array}[]{rl}E(e^{-sZ_{j,n+1}})=&E(e^{-s(\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1})}+1-e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}})\vskip 5.69054pt\\ =&E(e^{-s\sum_{k=1}^{N}a_{j,k}Z_{k,n}})S_{j}(s)A_{j}(-s)+1-U_{j,n}^{-}(s),\end{array}$ (82)
where
$U_{j,n}^{-}(s)=E(e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}}).$
By assuming that for each $n$, $Z_{j,n}$, $j=1,\ldots,N,$ are independent, then $Z^{n}(s_{1},\ldots,s_{N}):=E(e^{-\sum_{k=1}^{N}Z_{k,n}s_{k}})=\prod_{k=1}^{N}Z^{n}_{k}(s_{k})$. Now, by letting $n\to\infty$, (82) is rewritten for $j=1,\ldots,N$, $Re(s)=0$, as
$Z_{j}(s)=S_{j}(s)A_{j}(-s)\prod_{k=1}^{N}Z_{k}(a_{j,k}s)+1-U_{j}^{-}(s).$ (83)
Assume now that $A_{j,n}\sim exp(\lambda_{j})$, $j=1,\ldots,N$. Then, (83) is written for $Re(s)=0$ as
$(\lambda_{j}-s)Z_{j}(s)-\lambda_{j}S_{j}(s)\prod_{k=1}^{N}Z_{k}(a_{j,k}s)=(\lambda_{j}-s)(1-U_{j}^{-}(s)).$ (84)
Then,
  • The left-hand side of (84) is analytic in $Re(s)>0$, continuous in $Re(s)\geq 0$, and is also bounded.
  • The right-hand side of (84) is analytic in $Re(s)<0$, continuous in $Re(s)\leq 0$, and is also bounded.
Liouville’s theorem [30] states that for $Re(s)\geq 0$, $j=1,\ldots,N,$
$(\lambda_{j}-s)Z_{j}(s)-\lambda_{j}S_{j}(s)\prod_{k=1}^{N}Z_{k}(a_{j,k}s)=c_{0,j}+sc_{1,j}.$ (85)
For $s=0$, since $Z_{j}(0)=1$, (85) implies that $c_{0,j}=0$, $j=1,\ldots,N$. Moreover, setting $s=\lambda_{j}$ in (85) we found that
$c_{1,j}=-S_{j}(\lambda_{j})\prod_{l_{1}=1}^{N}Z_{l_{1}}(a_{j,k}\lambda_{j}),\,j=1,\ldots,N,$ (86)
and
$Z_{j}(s)=K_{j}(s)\prod_{l_{1}=1}^{N}Z_{l_{1}}(a_{j,l_{1}}s)+L_{j}(s).$ (87)
where
$K_{j}(s):=\frac{\lambda_{j}}{\lambda_{j}-s}S_{j}(s),\,L_{j}(s):=-\frac{s}{\lambda_{j}-s}c_{1,j}.$
In matrix notation (i.e., the vector of the marginal LSTs), (87) is written as
$\tilde{Z}(s)=K(s)\sum_{i=1}^{N}\prod_{j=1}^{N}\circ P_{i,j}\tilde{Z}(a_{i,j}s)+\tilde{L}(s),$ (88)
where,
$\prod_{j=1}^{N}\circ P_{i,j}\tilde{Z}(a_{i,j}s):=P_{i,1}\tilde{Z}(a_{i,1}s)\circ P_{i,2}\tilde{Z}(a_{i,2}s)\circ\ldots\circ P_{i,N}\tilde{Z}(a_{i,N}s),$ (89)
with $\circ$” denotes the Hadamard product, $P_{i,j}$ be an $N\times N$ matrix with the $(i,j)$ element is equal to 1, and all others equal to 0, and $K(s):=diag(K_{1}(s),\ldots,K_{N}(s))$, $L(s):=diag(L_{1}(s),\ldots,L_{N}(s))$. Note that (88), although seems to has a similar form as the one in (34), it is far more complicated since in any iterating step we have to appropriately substitute for any $\tilde{Z}(a_{i,j}s)$, $i,j=1,\ldots,N$, thus rapidly increasing the number of terms.
Remark 24
Consider the special case where $A=diag(a_{1},\ldots,a_{N})$. Then,
$Z_{j,n+1}=[a_{j}Z_{j,n}+S_{j,n}-A_{j,n+1}]^{+},\,j=1,\ldots,N,$ (90)
which corresponds to the following functional equation for the marginal LST
$Z_{j}(s)=Z_{j}(a_{j}s)S_{j}(s)A_{j}(-s)+1-U_{j}^{-}(s).$ (91)
Assuming that $A_{j,n}\sim exp(\lambda_{j})$, so that $A_{j}(s)=\frac{\lambda_{j}}{\lambda_{j}+s}$ $j=1,\ldots,N$, we have that for $Re(s)=0$,
$(\lambda_{j}-s)Z_{j}(s)-Z_{j}(a_{j}s)S_{j}(s)\lambda_{j}=(\lambda_{j}-s)(1-U_{j}^{-}(s)).$
Liouville’s theorem [30] implies that for $Re(s)\geq 0$,
$(\lambda_{j}-s)Z_{j}(s)-Z_{j}(a_{j}s)S_{j}(s)\lambda_{j}=c_{0,j}+c_{1,j}s.$
For $s=0$, we have that $c_{0,j}=0$, $j=1,\ldots,N$. Thus, for $Re(s)\geq 0$,
$Z_{j}(s)=Z_{j}(a_{j}s)S_{j}(s)\frac{\lambda_{j}}{\lambda_{j}-s}-\frac{sc_{1,j}}{\lambda_{j}-s}$ (92)
or equivalently, in matrix notation
$\tilde{Z}(s)=[S(s)\Lambda\circ I]\sum_{k=1}^{N}\tilde{P}^{(k)}\tilde{Z}(a_{k}s)+\tilde{V}(s),$ (93)
where $\circ"$ denotes the Hadamard product, $S(s)=diag(S_{1}(s),\ldots,S_{N}(s))$, $\Lambda=diag(\lambda_{1},\ldots,\lambda_{N})$, $\tilde{V}(s):=(I-A(-s))\tilde{c}$, $A(s):=diag(A_{1}(s),\ldots,A_{N}(s))$, $\tilde{c}=(c_{1,1},\ldots,c_{1,N})^{T}$. Note that (93) has the same form as (34), where $H(s)$ is equal now to $S(s)\Lambda\circ I$.
9.2 The joint Laplace-Stieltjes transform
Note that in order to come up with the expression in (83), we assumed that for each $n$, $Z_{j,n}$, $j=1,\ldots,N$ are independent. That assumption, although it creates a somehow more complicated vector valued functional equation in (88) (see also (93)), it can be treated similarly as those in the previous sections. In the following, we consider the joint LST of the stationary vector $(Z_{1},Z_{2})^{T}$, i.e., $f(s_{1},s_{2}):=E(e^{-s_{1}Z_{1}-s_{2}Z_{2}})$, and further drop the assumption of independence, but to keep the paper readable, we assume that $N=2$. Then,
$\begin{array}[]{rl}f(s_{1},s_{2}):=&E(e^{-s_{1}Z_{1}-s_{2}Z_{2}})=E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{+}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}})\vskip 5.69054pt\\ =&E(\left(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}+1-e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}}\right)\vskip 5.69054pt\\ &\times\left(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]}+1-e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}}\right))\vskip 5.69054pt\\ =&E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})+E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})\vskip 5.69054pt\\ &-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})+E(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &+1-E(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}})+E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}}).\end{array}$ (94)
Now,
$\begin{array}[]{rl}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})=&E(e^{-Z_{1}\sum_{j=1}^{2}a_{j,1}s_{j}-Z_{2}\sum_{j=1}^{2}a_{j,2}s_{j}})\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}\vskip 5.69054pt\\ =&\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j}),\vskip 5.69054pt\\ E(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]})=&\frac{\lambda_{k}S_{k}(s_{k})}{\lambda_{k}-s}f(a_{k,1}s_{k},a_{k,2}s_{k}),\,k=1,2.\end{array}$
Note that for $k=1,2,$
$\begin{array}[]{rl}e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]^{-}}=&\int_{y=0}^{\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}}\lambda_{k}e^{-\lambda_{k}y}dy+\int_{y=\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}}^{\infty}e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-y]}\lambda_{k}e^{-\lambda_{k}y}dy\vskip 5.69054pt\\ =&1-e^{-\lambda_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}]}+\lambda_{k}e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}]}\int_{y=\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}}^{\infty}e^{-(\lambda_{k}-s_{k})y}dy\vskip 5.69054pt\\ =&1+\frac{s_{k}}{\lambda_{k}-s_{k}}e^{-\lambda_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}]}.\end{array}$
Therefore,
$E(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]^{-}})=1+\frac{s_{k}S_{k}(\lambda_{k})}{\lambda_{k}-s_{k}}f(\lambda_{k}a_{k,1},\lambda_{k}a_{k,2}),\,k=1,2,$
and
$\begin{array}[]{r}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})=E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}))\vskip 5.69054pt\\ =E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})+\frac{s_{2}}{\lambda_{2}-s_{2}}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{1}S_{1}(s_{1})}{\lambda_{1}-s_{1}}f(a_{1,1}s_{1},a_{1,2}s_{1})+\frac{\lambda_{1}s_{2}S_{2}(\lambda_{2})S_{1}(s_{1})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2}),\end{array}$
and similarly,
$\begin{array}[]{c}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{2}S_{2}(s_{2})}{\lambda_{2}-s_{2}}f(a_{2,1}s_{2},a_{2,2}s_{2})+\frac{\lambda_{2}s_{1}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2}).\end{array}$
Moreover,
$\begin{array}[]{l}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})\vskip 5.69054pt\\ =E(\left(1+\frac{s_{1}}{\lambda_{1}-s_{1}}e^{-\lambda_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}]}\right)\left(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}\right))\vskip 5.69054pt\\ =1+\frac{s_{2}S_{2}(\lambda_{2})}{\lambda_{2}-s_{2}}f(\lambda_{2}a_{2,1},\lambda_{2}a_{2,2})+\frac{s_{1}S_{1}(\lambda_{1})}{\lambda_{1}-s_{1}}f(\lambda_{1}a_{1,1},\lambda_{1}a_{1,2})+\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+\lambda_{2}a_{2,1},\lambda_{1}a_{1,2}+\lambda_{2}a_{2,2}).\end{array}$
Substituting the above terms in (94) we come up with the following functional equation:
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j})-c_{2}(s_{1},s_{2})f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2})\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2})+c_{4}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}\lambda_{j},\sum_{j=1}^{2}a_{j,2}\lambda_{j}),\end{array}$ (95)
where
$\begin{array}[]{rl}c_{1}(s_{1},s_{2})=\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{2}(s_{1},s_{2})=\frac{\lambda_{1}s_{2}S_{1}(s_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},\vskip 5.69054pt\\ c_{3}(s_{1},s_{2})=\frac{s_{1}\lambda_{2}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{4}(s_{1},s_{2})=\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}.\end{array}$
[…]
Thus, as a first step in solving (95), we assume in the following that $A=diag(a_{1},a_{2})$ (where for notational convenience we denote $a_{1}$ for $a_{1,1}$, and $a_{2}$ for $a_{2,2}$). Note that for $A=diag(a_{1},a_{2})$, (95) is now written as:
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(a_{1}s_{1},a_{2}s_{2})-c_{2}(s_{1},s_{2})f(s_{1}a_{1},\lambda_{2}a_{2})-c_{3}(s_{1},s_{2})f(\lambda_{1}a_{1},s_{2}a_{2})+c_{4}(s_{1},s_{2})A_{1,1},\end{array}$ (96)
where from hereon we denote $A_{i,j}:=f(a_{1}^{i}\lambda_{1},a_{2}^{j}\lambda_{2})$, $i,j=1,2,\ldots$. Then, iterating (96) n1n-1 times results in
$\begin{array}[]{rl}f(s_{1},s_{2})=&C_{1,1}^{(n)}(s_{1},s_{2})A_{1,1}+\sum_{i=2}^{n}C_{i,1}^{(n)}(s_{1},s_{2})A_{i,1}+\sum_{j=2}^{n}C_{1,j}^{(n)}(s_{1},s_{2})A_{1,j}-\sum_{j=1}^{n}D_{j}^{(n)}(s_{1},s_{2})f(a_{1}^{n}s_{1},a_{2}^{j}\lambda_{2})\vskip 5.69054pt\\ &-\sum_{i=1}^{n}K_{i}^{(n)}(s_{1},s_{2})f(a_{1}^{i}\lambda_{1},a_{2}^{n}s_{2})+\prod_{k=0}^{n-1}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))f(a_{1}^{n}s_{1},a_{2}^{n}s_{2}),\end{array}$ (97)
where, $\sigma_{i,j}^{(k)}(s_{1},s_{2})=\sigma_{i,j}^{(k-1)}(\sigma^{(1)}_{i,j}(s_{1},s_{2}))$ (i.e., $\sigma^{(l)}(.)$ is the $l$th composition of $\sigma(.)$ with itself), with
$\sigma^{(1)}_{i,j}(s_{1},s_{2})=\left\{\begin{array}[]{rl}(a_{1}s_{1},a_{2}s_{2}),&i=j=1,\\ (a_{1}s_{1},a_{2}\lambda_{2}),&i=1,j=0,\\ (a_{1}\lambda_{1},a_{2}s_{2}),&i=0,j=1,\end{array}\right$.
and $\sigma^{(0)}_{1,1}(s_{1},s_{2}):=(s_{1},s_{2})$ (Note that $\sigma^{(1)}_{1,1}(s_{1},s_{2}):=A^{T}\tilde{s}$). Moreover, $C_{1,1}^{(n)}(s_{1},s_{2})$, $n=1,2,\ldots$ are recursively computed with $C_{1,1}^{(1)}(s_{1},s_{2}):=c_{4}(s_{1},s_{2})$, and
$\begin{array}[]{rl}C_{1,1}^{(n)}(s_{1},s_{2})=&C_{1,1}^{(n-1)}(s_{1},s_{2})+\prod_{k=0}^{n-2}c_{1}(\sigma^{(k)}_{1,1}(s_{1},s_{2}))c_{4}(\sigma^{(n-1)}_{1,1}(s_{1},s_{2}))-\sum_{j=1}^{n-1}D_{j}^{(n-1)}(s_{1},s_{2})c_{4}(a_{1}^{n-1}s_{1},a_{2}^{j}\lambda_{2})\vskip 5.69054pt\\ &-\sum_{i=1}^{n-1}K_{i}^{(n-1)}(s_{1},s_{2})c_{4}(a_{1}^{i-1}\lambda_{1},a_{2}^{n-1}s_{2}),\,n=2,3,\ldots.\end{array}$
and,
$\begin{array}[]{rl}C_{1,j}^{(n)}(s_{1},s_{2})=&\sum_{k=j-1}^{n-1}D_{j-1}^{(k)}(s_{1},s_{2})c_{3}(a_{1}^{k}s_{1},a_{2}^{j-1}\lambda_{2}),\,j=2,\ldots,n,\vskip 5.69054pt\\ C_{i,1}^{(n)}(s_{1},s_{2})=&\sum_{k=i-1}^{n-1}K_{i-1}^{(k)}(s_{1},s_{2})c_{2}(a_{1}^{i-1}\lambda_{1},a_{2}^{k}s_{2}),\,i=2,\ldots,n,\end{array}$
Furthermore $D_{1}^{(1)}(s_{1},s_{2}):=c_{2}(s_{1},s_{2})$ and for $n=2,3,\ldots,$
$\begin{array}[]{rl}D_{1}^{(n)}(s_{1},s_{2})=&\prod_{k=0}^{n-1}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))D_{1}^{(1)}(\sigma_{1,1}^{(n-1)}(s_{1},s_{2}))-\sum_{j=1}^{n-1}D_{j}^{(n-1)}(s_{1},s_{2})D_{1}^{(1)}(a_{1}^{n-1}s_{1},a_{2}^{j}\lambda_{2}),\vskip 5.69054pt\\ D_{j}^{(n)}(s_{1},s_{2})=&D_{j-1}^{(n-1)}(s_{1},s_{2})c_{1}(a_{1}^{n-1}s_{1},a_{2}^{j-1}\lambda_{2}),\,j=2,\ldots,n.\end{array}$
Similarly, $K_{1}^{(1)}(s_{1},s_{2}):=c_{3}(s_{1},s_{2})$ and for $n=2,3,\ldots,$
$\begin{array}[]{rl}K_{1}^{(n)}(s_{1},s_{2})=&\prod_{k=0}^{n-1}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))K_{1}^{(1)}(\sigma_{1,1}^{(n-1)}(s_{1},s_{2}))-\sum_{i=1}^{n-1}K_{i}^{(n-1)}(s_{1},s_{2})K_{1}^{(1)}(a_{1}^{i}\lambda_{1},a_{2}^{n-1}s_{2}),\vskip 5.69054pt\\ K_{i}^{(n)}(s_{1},s_{2})=&K_{i-1}^{(n-1)}(s_{1},s_{2})c_{1}(a_{1}^{i-1}\lambda_{1},a_{2}^{n-1}s_{2}),\,i=2,\ldots,n,\end{array}$
Letting $n\to\infty$, assuming $D_{j}^{(n)}(s_{1},s_{2})\to\tilde{D}_{j}(s_{1},s_{2})$, $K_{i}^{(n)}(s_{1},s_{2})\to\tilde{K}_{i}(s_{1},s_{2})$, $C_{1,1}^{(n)}(s_{1},s_{2})\to\tilde{C}_{1,1}(s_{1},s_{2})$, $C_{1,j}^{(n)}(s_{1},s_{2})\to\tilde{C}_{1,j}(s_{1},s_{2})$, $C_{i,1}^{(n)}(s_{1},s_{2})\to\tilde{C}_{i,1}(s_{1},s_{2})$, and having in mind that $f(0,0)=1$, (97) becomes
$\begin{array}[]{rl}f(s_{1},s_{2})=&\tilde{C}_{1,1}(s_{1},s_{2})A_{1,1}+\sum_{i=2}^{\infty}\tilde{C}_{i,1}(s_{1},s_{2})A_{i,1}+\sum_{j=2}^{\infty}\tilde{C}_{1,j}^{(n)}(s_{1},s_{2})A_{1,j}\vskip 5.69054pt\\ &-\sum_{j=1}^{\infty}\tilde{D}_{j}(s_{1},s_{2})f(0,a_{2}^{j}\lambda_{2})-\sum_{i=1}^{\infty}\tilde{K}_{i}(s_{1},s_{2})f(a_{1}^{i}\lambda_{1},0)+\prod_{k=0}^{\infty}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2})).\end{array}$ (98)
Thus, we have to find $A_{i,j}$, and $f(0,a_{2}^{j}\lambda_{2})$, $f(a_{1}^{i}\lambda_{1},0)$, $i,j=1,2,\ldots$. Note that as the number of iterations increases, i.e., $i\to\infty$, $A_{i,1}:=f(a_{1}^{i}\lambda_{1},a_{2}\lambda_{2})\to f(0,a_{2}\lambda_{2})$, and $j\to\infty$, $A_{1,j}:=f(a_{1}\lambda_{1},a_{2}^{j}\lambda_{2})\to f(a_{1}\lambda_{1},0)$ thus in practice, we need to obtain a few terms $A_{i,1}$, $A_{1,j}$, $i,j\geq 1$, and after that, all the terms converge to $f(0,a_{2}\lambda_{2})$, $f(a_{1}\lambda_{1},0)$.
We proceed first by obtaining $f(a_{1}^{i}\lambda_{1},0)$, $i=1,2,\ldots$. Set $s_{2}=0$ in (96) to obtain (note that $c_{2}(s_{1},0)=c_{4}(s_{1},0)=0$):
$f(s_{1},0)=c_{1}(s_{1},0)f(a_{1}s_{1},0)-c_{3}(s_{1},0)f(\lambda_{1}a_{1},0).$ (99)
It is easily realized that (99) is similar to the stationary version of a reflected autoregressive process analyzed in [12, Section 2.2]. So, by iterating (99) we have
$f(s_{1},0)=-f(\lambda_{1}a_{1},0)\sum_{j=0}^{\infty}c_{3}(a_{1}^{j}s_{1},0)\prod_{k=0}^{j-1}c_{1}(a_{1}^{k}s_{1},0)+\prod_{k=0}^{\infty}c_{1}(a_{1}^{k}s_{1},0).$ (100)
Substituting $s_{1}=a_{1}\lambda_{1}$ in (100) we obtain after some algebra
$f(\lambda_{1}a_{1},0)=\frac{\prod_{k=0}^{\infty}\frac{S_{1}(a_{1}^{k+1}\lambda_{1})}{1-a_{1}^{k+1}}}{1+\sum_{n=0}^{\infty}\frac{a_{1}^{n+1}S_{1}(\lambda_{1})}{1-a_{1}^{n+1}}\prod_{k=0}^{n-1}\frac{S_{1}(a_{1}^{k+1}\lambda_{1})}{1-a_{1}^{k+1}}}.$ (101)
Now, in (101) set $a_{1}\lambda_{1}$ instead of $\lambda_{1}$ to get:
$f(\lambda_{1}a_{1}^{2},0)=\frac{\prod_{k=0}^{\infty}\frac{S_{1}(a_{1}^{k+2}\lambda_{1})}{1-a_{1}^{k+1}}}{1+\sum_{n=0}^{\infty}\frac{a_{1}^{n+1}S_{1}(a_{1}\lambda_{1})}{1-a_{1}^{n+1}}\prod_{k=0}^{n-1}\frac{S_{1}(a_{1}^{k+2}\lambda_{1})}{1-a_{1}^{k+1}}},$
and continuing similarly,
$f(\lambda_{1}a_{1}^{i},0):=P_{i}=\frac{\prod_{k=0}^{\infty}\frac{S_{1}(a_{1}^{k+i}\lambda_{1})}{1-a_{1}^{k+1}}}{1+\sum_{n=0}^{\infty}\frac{a_{1}^{n+1}S_{1}(a_{1}^{i-1}\lambda_{1})}{1-a_{1}^{n+1}}\prod_{k=0}^{n-1}\frac{S_{1}(a_{1}^{k+i}\lambda_{1})}{1-a_{1}^{k+1}}},\,i=1,2,\ldots.$ (102)
By repeating the above procedure by letting $s_{1}=0$ in (99), we can obtain
$f(0,\lambda_{2}a_{2})=\frac{\prod_{k=0}^{\infty}\frac{S_{2}(a_{2}^{k+1}\lambda_{2})}{1-a_{2}^{k+1}}}{1+\sum_{n=0}^{\infty}\frac{a_{2}^{n+1}S_{2}(\lambda_{2})}{1-a_{2}^{n+1}}\prod_{k=0}^{n-1}\frac{S_{2}(a_{2}^{k+1}\lambda_{2})}{1-a_{2}^{k+1}}},$ (103)
and by iterating (103)
$f(0,\lambda_{2}a_{2}^{j}):=Q_{j}=\frac{\prod_{k=0}^{\infty}\frac{S_{2}(a_{2}^{k+j}\lambda_{2})}{1-a_{2}^{k+1}}}{1+\sum_{n=0}^{\infty}\frac{a_{2}^{n+j}S_{2}(a_{2}^{j-1}\lambda_{2})}{1-a_{2}^{n+1}}\prod_{k=0}^{n-1}\frac{S_{2}(a_{2}^{k+j}\lambda_{2})}{1-a_{2}^{k+1}}},\,j=1,2,\ldots.$ (104)
Remark 25
Note that by letting $i\to\infty$, the right-hand side of (102) converges to 1, which coincides with $\lim_{i\to\infty}f(\lambda_{1}a_{1}^{i},0)=f(0,0)=1$. Similarly, by letting $j\to\infty$, the right-hand side of (104) converges to 1, which coincides with $\lim_{j\to\infty}f(0,\lambda_{2}a_{2}^{j})=f(0,0)=1$.
Since $P_{i}$, $Q_{j}$, $i,j=1,2,\ldots,$ are known, we are ready to derive $A_{1,1}$, $A_{1,j}$ $j=2,\ldots$, $A_{i,1}$, $i=2,\ldots$ from (98). In particular, by substituting $(s_{1},s_{2})=(\lambda_{1}a_{1},\lambda_{2}a_{2})$, $(s_{1},s_{2})=(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})$, $n=2,3,\ldots$, $(s_{1},s_{2})=(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})$, $m=2,3,\ldots$ in (98), the unknown terms $A_{1,1}$, $A_{1,n}$ $n=2,\ldots$, $A_{m,1}$, $m=2,\ldots$ satisfy the following system of equations:
$\begin{array}[]{l}A_{1,1}(1-\tilde{C}_{1,1}(\lambda_{1}a_{1},\lambda_{2}a_{2}))-\sum_{i=2}^{\infty}\tilde{C}_{i,1}(\lambda_{1}a_{1},\lambda_{2}a_{2})A_{i,1}-\sum_{j=2}^{\infty}\tilde{C}_{1,j}(\lambda_{1}a_{1},\lambda_{2}a_{2})A_{1,j}\vskip 5.69054pt\\ =-\sum_{j=1}^{\infty}\tilde{D}_{j}(\lambda_{1}a_{1},\lambda_{2}a_{2})Q_{j}-\sum_{i=1}^{\infty}\tilde{K}_{i}(\lambda_{1}a_{1},\lambda_{2}a_{2})P_{i}+\prod_{k=0}^{\infty}c_{1}(\lambda_{1}a_{1}^{k+1},\lambda_{2}a_{2}^{k+1}),\vskip 5.69054pt\\ A_{1,n}(1-\tilde{C}_{1,n}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n}))-A_{1,1}\tilde{C}_{1,1}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})-\sum_{i=2}^{\infty}\tilde{C}_{i,1}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})A_{i,1}-\sum_{j=2,j\neq n}^{\infty}\tilde{C}_{1,j}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})A_{1,j}\vskip 5.69054pt\\ =-\sum_{j=1}^{\infty}\tilde{D}_{j}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})Q_{j}-\sum_{i=1}^{\infty}\tilde{K}_{i}(\lambda_{1}a_{1},\lambda_{2}a_{2}^{n})P_{i}+\prod_{k=0}^{\infty}c_{1}(\lambda_{1}a_{1}^{k+1},\lambda_{2}a_{2}^{k+n}),\,n=2,3,\ldots,\vskip 5.69054pt\\ A_{m,1}(1-\tilde{C}_{m,1}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2}))-A_{1,1}\tilde{C}_{1,1}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})-\sum_{i=2,i\neq m}^{\infty}\tilde{C}_{i,1}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})A_{i,1}-\sum_{j=2}^{\infty}\tilde{C}_{1,j}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})A_{1,j}\vskip 5.69054pt\\ =-\sum_{j=1}^{\infty}\tilde{D}_{j}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})Q_{j}-\sum_{i=1}^{\infty}\tilde{K}_{i}(\lambda_{1}a_{1}^{m},\lambda_{2}a_{2})P_{i}+\prod_{k=0}^{\infty}c_{1}(\lambda_{1}a_{1}^{k+m},\lambda_{2}a_{2}^{k+1}),\,m=2,3,\ldots\end{array}$
Having obtained $A_{1,1}$, $A_{i,1}$, $A_{1,j}$, $P_{i}$, $Q_{j}$, $i,j=1,2,\ldots,$ the joint LST $f(s_{1},s_{2})$ is given by (98). Note that in practice, since $a_{i}\in[0,1)$, $i=1,2,$ the series that appear above converge very rapidly, so we usually need a small finite number of iterations.
10 Conclusion and suggestions for future research
In this work, we dealt with vector-valued recursions between random vectors that lead to functional equations of the form (1), (2). In Section 2, we cope with the transient analysis of a Markov-modulated $M/G/1$-type reflected autoregressive process in which a vector-valued functional equation of the type (1) naturally arises. In Sections 3-5 we dealt with the stationary behavior of vector-valued recursions that arise from queueing processes and autoregressive processes where dependencies among random variables are dictated by (32), (39), and for which a similar vector-valued functional equation arises; see (2). In Section 6, we also cope with a queueing model, where dependencies are based on (32), but the vector-valued functional equation is somehow different compared to the others treated in this work, in the sense that the involved mappings $a_{i}(s)=s+\mu_{i}$, $i=1,\ldots,N$ are commutative, but there are not contractions. However, even in that case we were able to solve it iteratively. Section 7 is devoted to the analysis of a modulated 2-queue ASIP network with consumption, which led to an even more general functional equation. In Section 8, we focused on the modulated integer-valued reflected autoregressive process. Finally, in Section 9, we considered a reflected VAR(1) process, where the autoregressive parameter is now a given $N\times N$ matrix.
An interesting topic for future research is to cope with the case where we have vector-valued functional equations in which the involved contraction mappings are not commutative. For the scalar case (i.e., non-modulated), there exist some available scarce results, see [7], [29]. It would be interesting to further investigate what can still be accomplished both for the scalar and the vector-valued case, when we are dealing with a noncommutative contraction mapping. Other options for future research refer to the case where $R_{n}(X_{n})$ may take negative values.
Acknowledgements
I. Dimitriou gratefully acknowledges the Empirikion Foundation, Athens, Greece (www.empirikion.gr) for the financial support of this work. This work is dedicated to the memory of Antonis Kontis.